### v1

# Lipschitz-free spaces and Bossard’s reduction argument

Abstract.
Motivated by recent work of E. Basset, G. Lancien and A. Procházka, we use the reduction argument of Bossard to prove two results: if a separable Banach space is isomorphically universal for the class of Lipschitz-free spaces over the countable complete discrete metric spaces then it is isomorphically universal for the class of separable Banach spaces, and if a complete separable metric space is Lipschitz universal for the same class of metric spaces then it is Lipschitz universal for all separable metric spaces.
Key words and phrases:
Lipschitz-free space, universality, descriptive set theory
2020 Mathematics Subject Classification:
Primary 46B20; Secondary 03E15
1. Introduction
Let $M$ be a metric space with a base point denoted by 0. Let $\operatorname{Lip}_{0}(M)$ denote the Banach space of real-valued Lipschitz functions that vanish at 0, endowed with the norm that assigns to each $f\in\operatorname{Lip}_{0}(M)$ its optimal Lipschitz constant. Recall that the map ${\delta}:{M}\to{\operatorname{Lip}_{0}(M)^{*}}$ given by $\langle{f},{\delta(x)}\rangle=f(x)$, $x\in M$, $f\in\operatorname{Lip}_{0}(M)$, is an isometric embedding, and that we can define the Lipschitz-free space (hereafter free space) $\mathcal{F}({M})$ over $M$ as the closed linear span of the image $\delta(M)\subset\operatorname{Lip}_{0}(M)^{*}$ (where $\delta(M\setminus\{0\})$ is linearly independent).
For essential information on the fundamental properties of free spaces, we refer the reader to [godefroy:kalton:03, weaver:18]; in the latter they are known as Arens-Eells spaces. Free spaces have close ties to metric geometry, optimal transport theory and the linear and non-linear theory of Banach spaces, and have been the subject of intense study by the functional analysis community in the last two decades. Despite the ease with which they can be defined and the considerable progress made in understanding them over the intervening years, their structure is still relatively poorly understood and lots of “elementary” questions about them persist unanswered.
One of the enduring themes of this research has been to find conditions on metric spaces to decide whether their corresponding free spaces are linearly isomorphic or not. If $M$ is Lipschitz isomorphic to a subspace of $N$ then $\mathcal{F}({M})$ is linearly isomorphic to a subspace of $\mathcal{F}({N})$, and if $M$ and $N$ are Lipschitz isomorphic then $\mathcal{F}({M})$ and $\mathcal{F}({N})$ are linearly isomorphic. However, the converses to these statements fail dramatically. In particular, it is not easy to find families of metric spaces whose corresponding free spaces are pairwise non-isomorphic.
The first example of an uncountable family of separable metric spaces whose free spaces are pairwise non-isomorphic was given in [hajek:lancien:pernecka:16]. These metric spaces are all homeomorphic to the Cantor set $2^{\omega}$; the fact that their corresponding free spaces are pairwise non-isomorphic relies on the theorem of Johnson and Szankowski which states that there is no separable Banach space complementably universal for all separable Banach spaces.
Two more examples of such uncountable families are given in [basset:24, basset:lancien:prochazka:25]. In [basset:24], each metric space in the family is countable and complete, while in [basset:lancien:prochazka:25] each one is countable complete and discrete. The free spaces corresponding to the first family are non-isomorphic because the weak-fragmentability (or weak Szlenk) index of each one is distinct (and countable). The same holds in the case of the second family, this time with respect to the dentability index (not to be confused with the $w^{*}$-dentability index).
These results have consequences for universality. Recall that a metric space is purely 1-unrectifiable (p1u) if and only if it admits no Lipschitz isomorphic copy of any compact subset of $\mathbb{R}$ having strictly positive Lebesgue measure [kirchheim:94]. By [aliaga:gartand:petitjean:prochazka:22], a complete separable metric space $M$ is p1u if and only if $\mathcal{F}({M})$ has the Radon-Nikodým property (RNP); equivalently the dentability index of $\mathcal{F}({M})$ is countable. Because the dentability index of each free space associated with the second family is distinct, the set of indices must be unbounded in $\omega_{1}$. Thus, if a complete separable metric space has a free space that is isomorphically universal for the class of free spaces over countable complete discrete metric spaces, it cannot be p1u. Consequently, if a complete separable metric space is Lipschitz universal for the class of countable complete discrete metric spaces, it cannot be p1u [basset:lancien:prochazka:25]*Corollary 4.7.
I heard about the results in [basset:lancien:prochazka:25] in a talk given by E. Basset at the Banff International Research Station workshop on Functional and Metric Analysis and their Interactions, held at the Institute of Mathematics, University of Granada, Spain, in May 2025. It occurred to me that it might be possible to use descriptive set theory, specifically Bossard’s reduction argument (see [dodos:10]*p. 31), to build on their universality results. I would like to thank E. Basset, G. Lancien and A. Procházka for the fascinating exchange that ensued. I would also like to thank G. Godefroy for very interesting comments about the wider area. The two main results of the paper are given below. The second one was prompted by a question asked by G. Lancien.
Theorem 1.1.
Let $X$ be a separable Banach space that is isomorphically universal for the class of free spaces over countable complete discrete metric spaces. Then $X$ is isomorphically universal for the class of separable Banach spaces.
Theorem 1.2.
Let $M$ be a complete separable metric space that is Lipschitz universal for the class of countable complete discrete metric spaces. Then $M$ is Lipschitz universal for the class of separable metric spaces.
We prove some refinements of these results and make some related minor observations along the way. The application of descriptive set theory to Banach space theory can be traced to some seminal papers of Bourgain, e.g. [bourgain:80]. It was formalised later by Bossard in another seminal paper [bossard:02] and has since developed into a rich and beautiful subfield that boasts many powerful techniques and results, see e.g. [dodos:10] and references therein, and [cuth:doucha:kurka:22b] and related papers for more recent developments. We speculate that techniques from descriptive set theory could be used to prove more results about free spaces in the future. Finally, I should stress that I have left out most of the results in [basset:24, basset:lancien:prochazka:25] from this introduction purely in the interests of brevity, and I recommend that the interested reader should consult these papers in full.
2. Notation and preliminaries
We use standard Banach space notation throughout. The closed unit ball of a Banach space $X$ is denoted $B_{X}$. Given Banach spaces $X$ and $Y$, we write $X\hookrightarrow Y$ and $X\equiv Y$ when $X$ is linearly isomorphic to a subspace of $Y$ and $X$ is linearly isometric to $Y$, respectively. We also use these symbols in the context of metric spaces $M$ and $N$, with “Lipschitz” substituted for “linear”. By [godefroy:kalton:03]*Lemmas 2.2 and 2.3 (or [weaver:18]*Theorems 3.6 and 3.7), if $M\hookrightarrow N$ (in the Lipschitz sense) then $\mathcal{F}({M})\hookrightarrow\mathcal{F}({N})$ (in the linear sense). We say that the Banach space $X$ is isomorphically universal for a given class $\mathcal{C}$ of Banach spaces if $Y\hookrightarrow X$ whenever $Y\in\mathcal{C}$; Lipschitz universality of metric spaces is defined analogously.
For concepts pertaining to descriptive set theory, we follow [kechris:95]. The sets of natural numbers, countable ordinals and infinite sequences of natural numbers are denoted $\omega$, $\omega_{1}$ and $\omega^{\omega}$, respectively. We identify the power set of a set $\Gamma$ with $2^{\Gamma}$. We denote by $\mathbf{\Sigma}^{0}_{\alpha}$ and $\mathbf{\Pi}^{0}_{\alpha}$ the additive and multiplicative Borel classes of order $\alpha<\omega_{1}$ in a Polish space, so that $F_{\sigma}$, $G_{\delta}$ and $F_{\sigma\delta}$-sets are $\mathbf{\Sigma}^{0}_{2}$, $\mathbf{\Pi}^{0}_{2}$ and $\mathbf{\Pi}^{0}_{3}$-sets, and so on. Analytic and coanalytic subsets of Polish spaces are called $\mathbf{\Sigma}^{1}_{1}$ and $\mathbf{\Pi}^{1}_{1}$-sets, respectively. Given subsets $A\subset X$, $B\subset Y$ of Polish spaces $X$ and $Y$, we say that $A$ is reducible to $B$ if there is a continuous function ${f}:{X}\to{Y}$ such that $A=f^{-1}(B)$. If $\mathbf{\Gamma}(X)$ is a class of subsets of a Polish space $X$, a subset $B\subset Y$ of a Polish space $Y$ is called $\mathbf{\Gamma}$-complete if $B\in\mathbf{\Gamma}(Y)$ and $A$ is reducible to $B$ whenever $A\in\mathbf{\Gamma}(X)$ and $X$ is Polish and zero-dimensional.
The tree of finite (possibly empty) sequences of natural numbers is denoted $\omega^{<\omega}$. Given $s,t\in\omega^{<\omega}$ and $n\in\omega$, we define $|s|=\operatorname{dom}s\in\omega$ and $s^{\frown}n=s\cup\{(|s|,n)\}$, we write $s\preccurlyeq t$ when $t$ is an extension of $s$, i.e. $s=t\mathord{\upharpoonright}_{|s|}$, and $s\perp t$ when $s,t$ are incomparable, and finally we let $s\wedge t$ denote the greatest element of $\omega^{<\omega}$ less than both $s$ and $t$. The space of trees, i.e. downwards-closed subsets of $\omega^{<\omega}$, is denoted $\mathbf{Tr}$. With respect to the pointwise topology this is closed in $2^{\omega^{<\omega}}$ and hence compact. Within $\mathbf{Tr}$ we consider the subsets $\mathbf{WF}$ and $\mathbf{IF}$ of well-founded and ill-founded trees, which are well-known to be $\mathbf{\Pi}^{1}_{1}$ and $\mathbf{\Sigma}^{1}_{1}$-complete, respectively.
Next, we set up an appropriate descriptive set-theoretic framework for the class of separable infinite-dimensional free spaces. We shall consider the Polish space $0,\infty)^{\omega\times\omega}$ endowed with the pointwise topology, and the set $\mathcal{R}\subset[0,\infty)^{\omega\times\omega}$ of metrics on $\omega$. This space was introduced by Vershik in [vershik:98] and features again in e.g. [vershik:04] and [clemens:12] (in these works $\mathcal{R}$ is defined instead as the set of all pseudo-metrics on $\omega$, that is, functions $d\in[0,\infty)^{\omega\times\omega}$ that are metrics except that it is possible to have distinct $i,j\in\omega$ satisfying $d(i,j)=0$). It is easily seen that $\mathcal{R}$ is a $G_{\delta}$ in $0,\infty)^{\omega\times\omega}$ and is thus Polish as well. This space can be used to code the class of complete separable infinite metric spaces. Given such a space $(M,\rho)$, fix a dense sequence of distinct points $(x_{i})\subset M$ and consider $d\in\mathcal{R}$ given by $d(i,j)=\rho(x_{i},x_{j})$, $i,j\in\omega$; then $(M,\rho)\equiv M_{d}$, where $M_{d}$ is the completion of $(\omega,d)$. Conversely any such completion is separable and infinite. Because a free space is separable and infinite-dimensional if and only if its underlying metric space is separable and infinite, and because the free space over a metric space is isometric to that of its completion, $\mathcal{R}$ can be seen as a natural coding of the class of separable infinite-dimensional free spaces. As a change of base point does not alter the isometric structure of the corresponding free space, we shall always regard $0\in\omega$ as the base point.
Bossard coded the class of separable infinite-dimensional Banach spaces as a Borel subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure. The authors of [cuth:doucha:kurka:22a, cuth:doucha:kurka:22b] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$. First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ whose canonical extension to $c_{00}$ is a norm. Endowed with the pointwise topology, $\mathcal{B}$ is a $G_{\delta}$ in $0,\infty)^{V}$ and is thus Polish. Every separable and infinite-dimensional Banach space $(X,\left\|\cdot\right\|)$ has representatives in $\mathcal{B}$. Given a dense linearly independent sequence $(x_{i})\subset X$, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$. Then $X\equiv X_{\mu}$, where $X_{\mu}$ is the completion of $(V,\mu)$; conversely, every such completion is a separable and infinite-dimensional Banach space. We will work with $\mathcal{B}$ instead of $\mathcal{SB}$ because, in our opinion, it is easier to use and more compatible with $\mathcal{R}$. While not directly relevant to this paper, another advantage of using $\mathcal{B}$ is that Baire category arguments can be applied, e.g. the set of $\mu\in\mathcal{B}$ such that $X_{\mu}\equiv\mathbb{G}$, where $\mathbb{G}$ is the Gurariĭ space, is a dense $G_{\delta}$ [cuth:doucha:kurka:22b]*Theorem 4.1, hence $\mathbb{G}$ can be thought of as the generic separable Banach space. According to [cuth:doucha:kurka:22a]*Theorem 8, there is a Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ such that $\Theta(\mu)\equiv X_{\mu}$ for all $\mu\in\mathcal{B}$. It follows easily from this and [bossard:02]*Theorem 2.3 (ii) that, given any separable Banach space $X$, the set $\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow}:=\left\{{\mu\in\mathcal{B}}\;:\;\,{X_{\mu}\hookrightarrow X}\right\}$ is $\mathbf{\Sigma}^{1}_{1}$.
Concerning Polish versus Effros-Borel structures, we should mention that it is possible to code the class of complete infinite separable metric spaces using the set of closed infinite subsets of the Urysohn universal separable metric space $\mathbb{U}$, again equipped with the Effros-Borel structure, see e.g. [gao:09, gao:kechris:03]. However we shall work with Vershik’s space $\mathcal{R}$ for the reasons given above. Analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [vershik:04]*Theorems 1 and 3. Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [fonf:wojtaszczyk:08], could be regarded as the generic separable free space. A immediate consequence of this and [fonf:wojtaszczyk:08]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual. We mention this because Polish spaces of metrics have been used elsewhere to study approximation properties in free spaces [talimdjioski:23, smith:talimdjioski:24, talimdjioski:25].
3. The main results
Our first result will allow us to apply to $\mathcal{R}$ one of the observations about $\mathcal{B}$ recorded above.
Proposition 3.1.
There is a continuous map $\Xi$$\mathcal{R}$$\to{\mathcal{B}}$ such that $X_{\Xi(d)}\equiv\mathcal{F}({M_{d}})$ for all $d\in\mathcal{R}$.
Proof.
Recall that, regardless of $d\in\mathcal{R}$, the sequence $(\delta(i+1))_{i\in\omega}$ is always linearly independent and has dense span in $\mathcal{F}({M_{d}})$. Thus if we define $\Xi:\mathcal{R}\to\mathcal{B}$ by
$\Xi(d)(v)=\left\|{\sum_{i\in\omega}v(i)\delta(i+1)}\right\|_{\mathcal{F}({M_{d}})},\quad v\in V,$
we quickly conclude that $X_{\Xi(d)}\equiv\mathcal{F}({M_{d}})$.
The map $\Xi$ is continuous. Indeed, let $d\in\mathcal{R}$ and $v_{1},\ldots,v_{k}\in V$. There exists $N\in\omega$ such that $v_{\ell}(i)=0$ whenever $\ell=1,\ldots,k$ and $i\geqslant N$. Consider any $d^{\prime}\in\mathcal{R}$ such that $d^{\prime}(i,j)=d(i,j)$ whenever $i,j\leqslant N$. Then $\Xi(d^{\prime})(v_{\ell})=\Xi(d)(v_{\ell})$ for $\ell=1,\ldots,k$. This holds because, given $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$, by McShane’s extension theorem [weaver:18]*Theorem 1.33 there exists $g\in B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$, and vice-versa. Therefore $\Xi$ is continuous. ∎
Proposition 3.2.
Given a separable Banach space $X$ and a complete separable metric space $M$, the sets
$\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}:=\left\{{d\in\mathcal{R}}\;:\;\,{\mathcal{F}({M_{d}})\hookrightarrow X}\right\}\quad\text{and}\quad\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}:=\left\{{d\in\mathcal{R}}\;:\;\,{M_{d}\hookrightarrow M}\right\}$
are $\mathbf{\Sigma}^{1}_{1}$.
Proof.
Evidently $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\Xi^{-1}(\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$. Similarly to [clemens:12]*Lemma 4, the set
$\left\{{(d,d^{\prime})\in\mathcal{R}\times\mathcal{R}}\;:\;\,{M_{d}\hookrightarrow M_{d^{\prime}}}\right\}$
is $\mathbf{\Sigma}^{1}_{1}$, whence $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. ∎
Next, while not strictly necessary for what is to come, we determine the descriptive complexity of the classes of countably infinite complete discrete metric spaces and complete separable p1u metric spaces.
Proposition 3.3.
The set $\mathcal{R}_{\mathrm{dis}}:=\left\{{d\in\mathcal{R}}\;:\;\,{M_{d}\text{ is discrete}}\right\}$ is $\mathbf{\Pi}^{1}_{1}$.
Proof.
It is clear that $d\in\mathcal{R}_{\mathrm{dis}}$ if and only if every Cauchy sequence in $(\omega,d)$ is eventually constant. Given $k,m\in\omega$, define the sets
$A_{k,m}=\left\{{(d,(x_{i}))\in\mathcal{R}\times\omega^{\omega}}\;:\;\,{d(x_{i},x_{j})\leqslant 2^{-k}\text{ for all }i,j\geqslant m}\right\},$
and
$B_{m}=\left\{{(x_{i})\in\omega^{\omega}}\;:\;\,{x_{i}=x_{m}\text{ whenever }i\geqslant m}\right\},$
which are closed in $\mathcal{R}\times\omega^{\omega}$ and $\omega^{\omega}$, respectively. Then $(d,(x_{i}))\in A:=\bigcap_{k\in\omega}\bigcup_{m\in\omega}A_{k,m}$ if and only if $(x_{i})\in\omega^{\omega}$ is Cauchy in $(\omega,d)$, and $(x_{i})\in B:=\bigcup_{m\in\omega}B_{m}$ if and only if $(x_{i})$ is eventually constant. Therefore $\mathcal{R}\setminus\mathcal{R}_{\mathrm{dis}}$ is $\mathbf{\Sigma}^{1}_{1}$ because it is the image of $A\setminus(\mathcal{R}\times B)\subset\mathcal{R}\times\omega^{\omega}$ under the projection onto the first coordinate. ∎
Proposition 3.4.
[…]
Proof.
Using the Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ and [bossard:02]*Corollary 3.3, the set of $\mu\in\mathcal{B}$ such that $X_{\mu}$ has the RNP is $\mathbf{\Pi}^{1}_{1}$. Now apply Proposition 3.1 and [aliaga:gartand:petitjean:prochazka:22]*Theorem 4.6. ∎
Proposition 3.4 can be proved more directly using Kirchheim’s characterisation of p1u spaces mentioned above. Later, in Corollary 3.9, we upgrade $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$ to $\mathbf{\Pi}^{1}_{1}$-complete. For this and the main results we will need the following lemma, which is a tool for generating continuous maps from power sets to $\mathcal{R}$.
Lemma 3.5.
Let $(M,d)$ be a metric space, $A\subset\omega$, ${\phi_{0}}:{A}\to{M}$ and ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ injections with pairwise disjoint ranges, and let ${\tau}:{\omega\setminus A}\to{\Gamma}$, where $\Gamma$ is a set. Then the map ${\Psi}:{2^{\Gamma}}\to{M^{\omega}}$ given by
$\Psi(E)(n)=\begin{cases}\phi_{0}(n)&\text{if }n\in A,\\ \phi_{1}(n)&\text{if }n\in\omega\setminus A\text{ and }\tau(n)\notin E,\\ \phi_{2}(n)&\text{if }n\in\omega\setminus A\text{ and }\tau(n)\in E,\end{cases}\quad E\in 2^{\Gamma},$
is continuous. Moreover, the map ${\Delta}:{2^{\Gamma}}\to{[0,\infty)^{\omega\times\omega}}$ given by
$\Delta(E)(m,n)=d(\Psi(E)(m),\Psi(E)(n)),\quad m,n\in\omega,$
is continuous and takes values in $\mathcal{R}$. Finally, $(\omega,\Delta(E))$ is isometric to $(\Psi(E)(\omega),d)$.
Proof.
To check continuity of $\Psi$, let $E\in 2^{\Gamma}$ and let $F\subset\omega$ be finite. Define the finite sets $G=\tau(F\setminus A)\cap E$ and $H=\tau(F\setminus A)\setminus E$; then $G\subset E$, $H\cap E=\varnothing$ and $\Psi(E^{\prime})(n)=\Psi(E)(n)$ whenever $n\in F$, $E^{\prime}\in 2^{\Gamma}$, $G\subset E^{\prime}$ and $H\cap E^{\prime}=\varnothing$. Therefore $\Psi$ is continuous. As $\Delta$ is the composition of $\Psi$ and another continuous map, it is also continuous. Since $\phi_{0},\phi_{1},\phi_{2}$ are injective and have pairwise disjoint ranges, $\Psi(E)$ is injective for all $E\in 2^{\Gamma}$. It follows that $\Delta(E)\in\mathcal{R}$. The last assertion is obvious. ∎
Remark 3.6.
It is a simple exercise to show that $\mathcal{R}_{\mathrm{cpt}}:=\left\{{d\in\mathcal{R}}\;:\;\,{M_{d}\text{ is compact}}\right\}$ is $\mathbf{\Pi}^{0}_{3}$. As a straightforward application, we can show that $\mathcal{R}_{\mathrm{cpt}}$ is $\mathbf{\Pi}^{0}_{3}$-complete using Lemma 3.5. Indeed, let
$P=\left\{{E\in 2^{\omega\times\omega}}\;:\;\,{\left\{{j\in\omega}\;:\;\,{(i,j)\in E}\right\}\text{ is finite for all }i\in\omega}\right\}.$
According to [kechris:95]*Exercise 23.1, $P$ is $\mathbf{\Pi}^{0}_{3}$-complete. Now set $\Gamma=\omega\times\omega$, $M=\ell_{1}(\Gamma\cup\{\infty\})$, $A=\{0\}$ and $\phi_{0}=\{(0,0)\}$. Let $d$ be the usual metric on $M$, let ${\tau}:{\omega\setminus A}\to{\Gamma}$ be a bijection, let $(e_{\gamma})_{\gamma\in\Gamma\cup\{\infty\}}$ be the usual basis, and define injections ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ by
$\phi_{1}(n)=2^{-n}e_{\infty}\quad\text{and}\quad\phi_{2}(n)=2^{-p(\tau(n))}e_{\tau(n)},$
where ${p}:{\omega\times\omega}\to{\omega}$ is the projection onto the first coordinate. Evidently the ranges of $\phi_{0},\phi_{1},\phi_{2}$ are pairwise disjoint. Let $\Delta$ be the map furnished by Lemma 3.5. Then it is straightforward to check that $M_{\Delta(E)}=(\omega,\Delta(E))$ always, and this space is compact whenever $E\in P$ and not compact otherwise. Therefore $\Delta$ is a continuous reduction of $P$ to $\mathcal{R}_{\mathrm{cpt}}$, which implies that the latter is $\mathbf{\Pi}^{0}_{3}$-complete as claimed. The precise complexity of a few other standard classes of complete separable metric spaces can be determined in a similar manner.
We will be interested in metrics $\rho$ on $\omega^{<\omega}$ satisfying the following property.
Given $\varepsilon>0$ and an infinite branch $B\subset\omega^{<\omega}$, there exists $n\in\omega$ such that
$\rho(s,t)<\varepsilon$ whenever $s,t\in B$ and $|s|,|t|\geqslant n$. (3.1)
Lemma 3.7.
Let $(M,\theta)$ be a separable metric space, let $(x_{i})$ be a dense sequence in $M$ whose elements repeat infinitely often, and let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.1). Then the metric $d$ on $\omega^{<\omega}$ defined by
$d(s,t)=\rho(s,t)+\theta(x_{|s|},x_{|t|}),$
has the property that $(M,\theta)$ embeds isometrically into the completion $\overline{{(B,d)}}$ of $(B,d)$ whenever $B\subset\omega^{<\omega}$ is an infinite branch.
Proof.
Let $B\subset\omega^{<\omega}$ be an infinite branch. Then there exists $q\in\omega^{\omega}$ such that $B=\left\{{q\mathord{\upharpoonright}_{n}}\;:\;\,{n\in\omega}\right\}$. We define an isometric embedding ${\phi}:{M}\to{\overline{{(B,d)}}}$ as follows. Given $x\in M$, because the elements of $(x_{i})$ repeat infinitely often, there exist $i_{0}<i_{1}<i_{2}<\ldots$ such that $\theta(x_{i_{k}},x)\to 0$. Given that $\rho$ satisfies (3.1), $(q\mathord{\upharpoonright}_{i_{k}})$ is $\rho$-Cauchy, hence $(q\mathord{\upharpoonright}_{i_{k}})$ is also $d$-Cauchy. Let $\phi(x)=\lim_{k}q\mathord{\upharpoonright}_{i_{k}}\in\overline{{(B,d)}}$. This is well-defined because if $j_{0}<j_{1}<j_{2}<\ldots$ also satisfy $\theta(x_{j_{k}},x)\to 0$ then $d(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})\to 0$, again by (3.1). Finally, we check that $\phi$ is an isometric embedding. With $x$ and the $i_{k}$ as above, let $y\in M$ and this time let $j_{0}<j_{1}<j_{2}<\ldots$ such that $\theta(y,x_{j_{k}})\to 0$. Then once more by (3.1),
$d(\phi(x),\phi(y))=\lim_{k}d(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\lim_{k}\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})=\theta(x,y).\qed$
If $(M,\theta)$ above is perfect then there is no need to insist that the elements of $(x_{i})$ repeat infinitely often.
We now define a metric $\rho$ on $\omega^{<\omega}$ that we fix hereafter. Let $(e_{s})_{s\in\omega^{<\omega}}$ denote the usual basis of $c_{00}(\omega^{<\omega})$ and define the injection ${\Omega}:{\omega^{<\omega}}\to{c_{00}(\omega^{<\omega})}$ by
$\Omega(s)=\sum_{r\preccurlyeq s}2^{-|r|}e_{r},\quad s\in\omega^{<\omega}.$
Then define $\rho(s,t)=\left\|{\Omega(s)-\Omega(t)}\right\|_{\infty}$, $s,t\in\omega^{<\omega}$. We see that $\rho$ satisfies (3.1). If $s\perp t$ then $s\wedge t\prec s,t$ and
$\rho(s,t)=\left\|{\sum_{s\wedge t\prec r\preccurlyeq s}2^{-|r|}e_{r}-\sum_{s\wedge t\prec r\preccurlyeq t}2^{-|r|}e_{r}}\right\|_{\infty}=2^{-|s\wedge t|-1},$ (3.2)
and we get the same expression for $\rho(s,t)$ if $s$ and $t$ are distinct and comparable. Hence (3.1) holds.
Lemma 3.8.
Let $d$ be the metric from Lemma 3.7, where $(M,\theta)$ is arbitrary and $\rho$ is as above. Then $(\omega^{<\omega},d)$ is discrete, and $(T,d)$ is in addition complete whenever $T\in\mathbf{WF}$.
Proof.
Fix $s\in\omega^{<\omega}$. Given $t\in\omega^{<\omega}\setminus\{s\}$, by (3.2) we have
$d(s,t)\geqslant\rho(s,t)=2^{-|s\wedge t|-1}\geqslant 2^{-|s|-1}.$
Hence $d(s,t)\geqslant 2^{-|s|-1}$ whenever $t\in\omega^{<\omega}\setminus\{s\}$, meaning that $d$ is discrete.
Now let $T\in\mathbf{WF}$ and let $(s_{n})\subset T$ be $d$-Cauchy. We find a strictly increasing finite sequence $r_{0}\prec r_{1}\prec r_{2}\prec\ldots$ of elements of $\omega^{<\omega}$ and infinite subsets $L_{0}\supset L_{1}\supset L_{2}\supset\ldots$ of $\omega$ such that $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$. Begin by setting $r_{0}=\varnothing$ and $L_{0}=\omega$. Given $r_{k}$ and $L_{k}$, either there exists $m\in\omega$ and an infinite subset $L_{k+1}\subset L_{k}$ such that $r_{k}^{\frown}m\preccurlyeq s_{n}$ for all $n\in L_{k+1}$, or not. If the first case holds, set $r_{k+1}=r_{k}^{\frown}m$ and continue the process, else terminate it. Because $T\in\mathbf{WF}$, this process must terminate at some finite stage $k$. By construction we know that $L_{k}$ is infinite and $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$ but also, given $m\in\omega$, the set $H_{m}:=\left\{{n\in L_{k}}\;:\;\,{r_{k}^{\frown}m\preccurlyeq s_{n}}\right\}$ must be finite. If $n_{1}\in H_{m_{1}}$ and $n_{2}\in H_{m_{2}}$ for distinct $m_{1},m_{2}\in\omega$, then $s_{n_{1}}\perp s_{n_{2}}$ and $s_{n_{1}}\wedge s_{n_{2}}=r_{k}$, giving
$d(s_{n_{1}},s_{n_{2}})\geqslant\rho(s_{n_{1}},s_{n_{2}})=2^{-|r_{k}|-1}.$
Because $(s_{n})$ is $d$-Cauchy, it follows that $H_{m}$ is empty for all but finitely many $m\in\omega$, and thus $s_{n}=r_{k}$ for infinitely many $n\in L_{k}$. Again because $(s_{n})$ is $d$-Cauchy, we conclude that it must converge to $r_{k}$. This proves that $(T,d)$ is complete. ∎
The proofs of Theorem 1.1 and Theorem 1.2 use Bossard’s reduction argument [bossard:02]*Theorem 3.2.
Proof of Theorem 1.1.
Let $X$ be a separable Banach space. By Proposition 3.2, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Assume that $\mathcal{R}_{\mathrm{dis}}\subset\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. Let $\theta$ be the usual metric on $B_{c_{0}}$ inherited from $c_{0}$, and let $(\omega^{<\omega},d)$ be the metric space obtained from Lemma 3.7 applied to $(B_{c_{0}},\theta)$. It is evident that $d\leqslant\frac{5}{2}$. Let $M=\omega^{<\omega}\cup(\omega^{<\omega}\times\{0\})$ and extend $d$ to a metric on $M$ by setting $d(x,y)=\frac{5}{2}$ whenever $x\neq y$ and $\{x,y\}\setminus\omega^{<\omega}$ is non-empty (it is straightforward to check that the extended function satisfies the triangle inequality).
Set $A=\phi_{0}=\varnothing$, $\Gamma=\omega^{<\omega}$, let ${\tau}:{\omega}\to{\Gamma}$ be a bijection, and define injections ${\phi_{1},\phi_{2}}:{\omega}\to{M}$ by $\phi_{1}(n)=(\tau(n),0)$ and $\phi_{2}(n)=\tau(n)$, $n\in\omega$. Let ${\Delta}:{2^{\Gamma}}\to{\mathcal{R}}$ be the continuous map furnished by Lemma 3.5, which we then restrict to $\mathbf{Tr}$.
Let $T\in\mathbf{Tr}$. We observe that $\Psi(T)(\omega)=T\cup((\omega^{<\omega}\setminus T)\times\{0\})$. Using Lemma 3.8, $(T,d)$ is discrete and complete if $T\in\mathbf{WF}$. Hence $(\omega,\Delta(T))$, which is isometric to $(\Psi(T)(\omega),d)$, is also discrete and complete because every point in $(\omega^{<\omega}\setminus T)\times\{0\}$ is $d$-separated from everything else by $\frac{5}{2}$. Therefore $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$. Now let $T\in\mathbf{IF}$ and let $B\subset T$ be an infinite branch. Then the completion $M_{\Delta(T)}$ of $(\omega,\Delta(T))$ contains an isometric copy of the completion of $(B,d)$. By Lemma 3.7, $B_{c_{0}}\hookrightarrow M_{\Delta(T)}$.
As $\mathcal{R}_{\mathrm{dis}}\subset\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$, we have $\mathbf{WF}=\Delta^{-1}(\mathcal{R}_{\mathrm{dis}})\subset\Delta^{-1}(\langle X\rangle^{\mathcal{R}}_{\hookrightarrow})$. Since $\mathbf{WF}$ is $\mathbf{\Pi}^{1}_{1}$-complete and $\Delta^{-1}(\langle X\rangle^{\mathcal{R}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$, there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. Therefore $\mathcal{F}({B_{c_{0}}})\hookrightarrow\mathcal{F}({M_{d}})\hookrightarrow X$. Because $Y\hookrightarrow\mathcal{F}({B_{c_{0}}})$ whenever $Y$ is a separable Banach space ([dutrieux:ferenczi:06]*p. 1042 and [kaufmann:15]*Corollary 3.3), the conclusion follows. ∎
Proof of Theorem 1.2.
Let $M$ be a complete separable metric space. By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. By repeating part of the proof above, this time applying Lemma 3.7 to $\mathbb{U}$, we obtain a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ if $T\in\mathbf{WF}$ and $\mathbb{U}\hookrightarrow M_{\Delta(T)}$ otherwise. By the last part of the proof above, there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$, which implies $\mathbb{U}\hookrightarrow M_{\Delta(T)}\hookrightarrow M$. ∎
Corollary 3.9.
The sets $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$ are $\mathbf{\Pi}^{1}_{1}$-complete.
Proof.
The maps $\Delta$ in the proofs of Theorems 1.1 and 1.2 witness the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$. ∎
We finish by making a few remarks about Theorems 1.1 and 1.2. First, it is clear that $M_{\Delta(T)}$ in the proof of Theorem 1.1 is always bounded, so if we require that $X$ is universal for the class of free spaces over countable complete discrete spaces that are in addition bounded, the outcome is the same. Of course, in the proof of Theorem 1.2 we can replace $\mathbb{U}$ by any separable Lipschitz universal space such as $c_{0}$ (by Aharoni’s theorem, see e.g. [benyamini:lindenstrauss:00]*Theorem 7.11). If we add boundedness to the list of properties in Theorem 1.2 then $M$ is Lipschitz universal for all bounded separable metric spaces. This is because we can replace $\mathbb{U}$ in the proof e.g. by any ball in $\mathbb{U}$ (see [vershik:98]*p. 922) or any ball in $c_{0}$.
Finally, we can add the property of being uniformly concave to the lists above and obtain the same conclusions. According to [weaver:18]*Definition 3.33, $(M,d)$ is uniformly concave if, given distinct $x,z\in M$ and $\varepsilon>0$, there exists $\delta>0$ such that $d(x,z)\leqslant d(x,y)+d(y,z)-\delta$ whenever $d(x,y),d(y,z)\geqslant\varepsilon$. Uniformly concave metric spaces are relevant to the extremal structure of free spaces (see e.g. [weaver:18]*Theorem 3.39).
It can be shown that $(\omega,\Delta(T))=M_{\Delta(T)}$ is uniformly concave whenever $T\in\mathbf{WF}$. This holds because $(\omega^{<\omega},\rho)$ is uniformly concave (with $\delta$ independent of $\varepsilon$). To see this we require a quick lemma.
Lemma 3.10.
If $s,t,u\in\omega^{<\omega}$ and $s\wedge u\prec s\wedge t$ then $s\wedge u=t\wedge u$.
Proof.
As $s\wedge u\preccurlyeq u$ and $s\wedge u\prec s\wedge t\preccurlyeq t$, we have $s\wedge u\preccurlyeq t\wedge u$. Because $s\wedge t,t\wedge u\preccurlyeq t$, they are comparable. However, if $s\wedge t\preccurlyeq t\wedge u\preccurlyeq u$ then $s\wedge t\preccurlyeq s\wedge u$, which is not the case. Thus $t\wedge u\prec s\wedge t\preccurlyeq s$, giving $t\wedge u\preccurlyeq s\wedge u$. ∎
Now let $s,t,u\in\omega^{<\omega}$ be distinct and set $\delta=\min\{2^{-|s|-1},2^{-|u|-1}\}>0$. We will show that $\rho(s,u)\leqslant\rho(s,t)+\rho(t,u)-\delta$, thus demonstrating that $(\omega^{<\omega},\rho)$ is uniformly concave. If $|s\wedge t|\leqslant|s\wedge u|$ then
$\rho(s,t)+\rho(t,u)=2^{-|s\wedge t|-1}+2^{-|t\wedge u|-1}\geqslant 2^{-|s\wedge u|-1}+2^{-|u|-1}\geqslant\rho(s,u)+\delta;$
and we get the same result if $|t\wedge u|\leqslant|s\wedge u|$. On the other hand, if $|s\wedge u|<|s\wedge t|$ then $s\wedge u\prec s\wedge t$ because these elements are comparable, thus $s\wedge u=t\wedge u$ by Lemma 3.10. Then we estimate
$\rho(s,t)+\rho(t,u)=2^{-|s\wedge t|-1}+2^{-|t\wedge u|-1}\geqslant 2^{-|s|-1}+2^{|s\wedge u|-1}\geqslant\delta+\rho(s,u).$
We reach the same conclusion if $|s\wedge u|<|t\wedge u|$.

### v2

@@
 Abstract.
-Motivated by recent work of E.
-Basset, G.
-Lancien and A.
-Procházka, we use the reduction argument of Bossard to prove two results: if a separable Banach space is isomorphically universal for the class of Lipschitz-free spaces over the countable complete discrete metric spaces then it is isomorphically universal for the class of separable Banach spaces, and if a complete separable metric space is Lipschitz universal for the same class of metric spaces then it is Lipschitz universal for all separable metric spaces.
+We set up a descriptive set-theoretic framework to study Lipschitz-free spaces and use the reduction argument of Bossard to prove several results.
+We prove two universality results: if a separable Banach space is isomorphically universal for the class of Lipschitz-free spaces over the countable complete discrete metric spaces then it is isomorphically universal for the class of separable Banach spaces, and if a complete separable metric space is Lipschitz universal for the same class of metric spaces then it is Lipschitz universal for all separable metric spaces.
+We also show that there exist countable complete discrete metric spaces whose Lipschitz-free spaces fail the bounded approximation property and are thus not isomorphic to any dual Banach space.
+Finally, we calculate the descriptive complexity of the classes of separable Banach spaces and separable Lipschitz-free spaces having the approximation property.
 Key words and phrases:
-Lipschitz-free space, universality, descriptive set theory
+Lipschitz-free space, universality, descriptive set theory, approximation property, bounded approximation property
 2020 Mathematics Subject Classification:
@@
 Let $\operatorname{Lip}_{0}(M)$ denote the Banach space of real-valued Lipschitz functions that vanish at 0, endowed with the norm that assigns to each $f\in\operatorname{Lip}_{0}(M)$ its optimal Lipschitz constant.
-Recall that the map ${\delta}:{M}\to{\operatorname{Lip}_{0}(M)^{*}}$ given by $\langle{f},{\delta(x)}\rangle=f(x)$, $x\in M$, $f\in\operatorname{Lip}_{0}(M)$, is an isometric embedding, and that we can define the Lipschitz-free space (hereafter free space) $\mathcal{F}({M})$ over $M$ as the closed linear span of the image $\delta(M)\subset\operatorname{Lip}_{0}(M)^{*}$ (where $\delta(M\setminus\{0\})$ is linearly independent).
+Recall that the map ${\delta}:{M}\to{\operatorname{Lip}_{0}(M)^{*}}$ given by $\langle{f},{\delta(x)}\rangle=f(x)$, $x\in M$, $f\in\operatorname{Lip}_{0}(M)$, is an isometric embedding, and that we can define the Lipschitz-free space (hereafter free space) $\mathcal{F}({M})$ over $M$ as the closed linear span of the image $\delta(M)\subseteq\operatorname{Lip}_{0}(M)^{*}$ (where $\delta(M\setminus\{0\})$ is linearly independent).
 For essential information on the fundamental properties of free spaces, we refer the reader to [godefroy:kalton:03, weaver:18]; in the latter they are known as Arens-Eells spaces.
@@
 The first example of an uncountable family of separable metric spaces whose free spaces are pairwise non-isomorphic was given in [hajek:lancien:pernecka:16].
-These metric spaces are all homeomorphic to the Cantor set $2^{\omega}$; the fact that their corresponding free spaces are pairwise non-isomorphic relies on the theorem of Johnson and Szankowski which states that there is no separable Banach space complementably universal for all separable Banach spaces.
-Two more examples of such uncountable families are given in [basset:24, basset:lancien:prochazka:25].
+These metric spaces are all homeomorphic to the Cantor set $2^{\omega}$.
+Two more uncountable families of this kind are given in [basset:24, basset:lancien:prochazka:25].
 In [basset:24], each metric space in the family is countable and complete, while in [basset:lancien:prochazka:25] each one is countable complete and discrete.
@@
 Consequently, if a complete separable metric space is Lipschitz universal for the class of countable complete discrete metric spaces, it cannot be p1u [basset:lancien:prochazka:25]*Corollary 4.7.
+In this paper we use a tool from descriptive set theory, specifically Bossard’s reduction argument (see [dodos:10]*p. 31), to build on this universality result.
+Theorem 1.1.
+Let $X$ be a separable Banach space that is isomorphically universal for the class of free spaces over countable complete discrete metric spaces.
+Then $X$ is isomorphically universal for the class of separable Banach spaces.
+Theorem 1.2.
+Let $M$ be a complete separable metric space that is Lipschitz universal for the class of countable complete discrete metric spaces.
+Then $M$ is Lipschitz universal for the class of separable metric spaces.
+Another enduring strand of this research has been to identify which free spaces have or fail one of Grothendieck’s classical approximation properties.
+Recall that a Banach space $X$ is said to have the approximation property (AP) if, given compact $K\subseteq X$ and $\varepsilon>0$, there exists a bounded linear finite-rank operator ${T}:{X}\to{X}$ such that $\left\|{Tx-x}\right\|\leqslant\varepsilon$ whenever $x\in K$.
+If there exists $\lambda\geqslant 1$ such that the operator $T$ can always be chosen to satisfy $\left\|{T}\right\|\leqslant\lambda$, then $X$ is said to have the $\lambda$-bounded approximation property ($\lambda$-BAP).
+We say that $X$ has the bounded approximation property (BAP) if it has the $\lambda$-BAP for some $\lambda$, and the metric approximation property (MAP) if it has the $1$-BAP.
+The literature on free spaces and approximation properties is too abundant to be summarised in this short paper; we invite the interested reader to consult [godefroy:24] and the introduction of [smith:talimdjioski:23] and references therein for more information.
+We only need to recall two results for our purposes.
+First, it follows immediately from [godefroy:kalton:03]*Theorem 3.1 that if a separable Banach space $X$ fails the AP then so does $\mathcal{F}({X})$.
+Second, [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8 implies that $\mathcal{F}({M})$ has the AP whenever $M$ is countable discrete and complete.
+To the best of my knowledge, to date there has been no example of a complete separable p1u metric space whose free space has the AP but fails the BAP.
+The next result shows that there are such examples and, moreover, they can be chosen to be countable complete and discrete.
+Again, this result is based on Bossard’s argument.
+Another crucial ingredient is the fact that, relative to the established descriptive set-theoretic frameworks for separable Banach spaces (see Section 2), the class of separable Banach spaces that satisfies the BAP is Borel (though being analytic would have sufficed) [ghawadrah:17].
+Theorem 1.3.
+There exist countable complete discrete metric spaces $M$ such that $\mathcal{F}({M})$ fails the BAP. Consequently, $\mathcal{F}({M})$ is not isomorphic to a dual space.
+This result can be contrasted with [aliaga:gartand:petitjean:prochazka:22]*Theorem 3.2 which states, among other things, that $\mathcal{F}({M})$ is isometric to a dual space whenever $M$ is a p1u metric space that is proper, i.e.
+its closed balls are compact.
+It can also be regarded in the light of Kalton’s problem of whether $\mathcal{F}({M})$ has the BAP whenever $M$ is uniformly discrete; see [kalton:04]*p.
+185 and e.g.
+[godefroy:24]*Problem 4.1 for more context.
+In this vein, there exists a uniformly discrete space whose free space is not isometric to a dual space (though this particular example has the MAP) [garcia-lirola:petitjean:prochazka:ruedazoca:18]*Example 5.8.
+We prove the results above using Lemma 3.10, which is one of our two main tools.
+The other main tool, Lemma 3.9, will be needed for our final main result, Theorem 4.1, which shows that the classes of separable Banach spaces and separable free spaces having the AP are, relative to the appropriate descriptive set-theoretic interpretations, complete coanalytic sets.
+(We remark that this is not the first time Polish spaces of metrics have been used to study free spaces and approximation properties; see [talimdjioski:23, smith:talimdjioski:24, talimdjioski:25].)
+We prove these theorems in Sections 3 and 4, and make some related observations along the way.
+The application of descriptive set theory to Banach space theory can be traced to some seminal papers of Bourgain, e.g.
+[bourgain:80].
+It was formalised later by Bossard in another seminal paper [bossard:02] and has since developed into a rich and beautiful subfield that boasts many powerful techniques and results, see e.g.
+[dodos:10] and references therein, and [cuth:doucha:kurka:22b] and related papers for more recent developments.
+I speculate that techniques from descriptive set theory could be used to prove more results about free spaces in the future.
 I heard about the results in [basset:lancien:prochazka:25] in a talk given by E.
@@
 Lancien and A.
-Procházka for the fascinating exchange that ensued.
-I would also like to thank G.
-Godefroy for very interesting comments about the wider area.
-The two main results of the paper are given below.
-The second one was prompted by a question asked by G.
+Procházka for the fascinating exchanges that ensued.
+In particular, Theorem 1.2 was prompted by a question asked by G.
 Lancien.
-Theorem 1.1.
-Let $X$ be a separable Banach space that is isomorphically universal for the class of free spaces over countable complete discrete metric spaces.
-Then $X$ is isomorphically universal for the class of separable Banach spaces.
-Theorem 1.2.
-Let $M$ be a complete separable metric space that is Lipschitz universal for the class of countable complete discrete metric spaces.
-Then $M$ is Lipschitz universal for the class of separable metric spaces.
-We prove some refinements of these results and make some related minor observations along the way.
-The application of descriptive set theory to Banach space theory can be traced to some seminal papers of Bourgain, e.g.
-[bourgain:80].
-It was formalised later by Bossard in another seminal paper [bossard:02] and has since developed into a rich and beautiful subfield that boasts many powerful techniques and results, see e.g.
-[dodos:10] and references therein, and [cuth:doucha:kurka:22b] and related papers for more recent developments.
-We speculate that techniques from descriptive set theory could be used to prove more results about free spaces in the future.
+I would also like to thank M.
+Cúth, G.
+Godefroy and A.
+Rueda Zoca for several remarks which helped me when I was preparing this paper.
 Finally, I should stress that I have left out most of the results in [basset:24, basset:lancien:prochazka:25] from this introduction purely in the interests of brevity, and I recommend that the interested reader should consult these papers in full.
@@
 We also use these symbols in the context of metric spaces $M$ and $N$, with “Lipschitz” substituted for “linear”.
-By [godefroy:kalton:03]*Lemmas 2.2 and 2.3 (or [weaver:18]*Theorems 3.6 and 3.7), if $M\hookrightarrow N$ (in the Lipschitz sense) then $\mathcal{F}({M})\hookrightarrow\mathcal{F}({N})$ (in the linear sense).
+By [weaver:18]*Theorems 3.6 and 3.7, if $M\hookrightarrow N$ (in the Lipschitz sense) then $\mathcal{F}({M})\hookrightarrow\mathcal{F}({N})$ (in the linear sense).
 We say that the Banach space $X$ is isomorphically universal for a given class $\mathcal{C}$ of Banach spaces if $Y\hookrightarrow X$ whenever $Y\in\mathcal{C}$; Lipschitz universality of metric spaces is defined analogously.
@@
 We denote by $\mathbf{\Sigma}^{0}_{\alpha}$ and $\mathbf{\Pi}^{0}_{\alpha}$ the additive and multiplicative Borel classes of order $\alpha<\omega_{1}$ in a Polish space, so that $F_{\sigma}$, $G_{\delta}$ and $F_{\sigma\delta}$-sets are $\mathbf{\Sigma}^{0}_{2}$, $\mathbf{\Pi}^{0}_{2}$ and $\mathbf{\Pi}^{0}_{3}$-sets, and so on.
-Analytic and coanalytic subsets of Polish spaces are called $\mathbf{\Sigma}^{1}_{1}$ and $\mathbf{\Pi}^{1}_{1}$-sets, respectively.
-Given subsets $A\subset X$, $B\subset Y$ of Polish spaces $X$ and $Y$, we say that $A$ is reducible to $B$ if there is a continuous function ${f}:{X}\to{Y}$ such that $A=f^{-1}(B)$.
-If $\mathbf{\Gamma}(X)$ is a class of subsets of a Polish space $X$, a subset $B\subset Y$ of a Polish space $Y$ is called $\mathbf{\Gamma}$-complete if $B\in\mathbf{\Gamma}(Y)$ and $A$ is reducible to $B$ whenever $A\in\mathbf{\Gamma}(X)$ and $X$ is Polish and zero-dimensional.
+Analytic, coanalytic and Borel subsets of Polish spaces are called $\mathbf{\Sigma}^{1}_{1}$, $\mathbf{\Pi}^{1}_{1}$ and $\mathbf{\Delta}^{1}_{1}$-sets, respectively.
+Given subsets $A\subseteq X$, $B\subseteq Y$ of Polish spaces $X$ and $Y$, we say that $A$ is reducible to $B$ if there is a continuous function ${f}:{X}\to{Y}$ such that $A=f^{-1}(B)$.
+If $\mathbf{\Gamma}(X)$ is a class of subsets of a Polish space $X$, a subset $B\subseteq Y$ of a Polish space $Y$ is called $\mathbf{\Gamma}$-hard if $A$ is reducible to $B$ whenever $A\in\mathbf{\Gamma}(X)$ and $X$ is Polish and zero-dimensional, and $\mathbf{\Gamma}$-complete if $B$ is $\mathbf{\Gamma}$-hard and in addition belongs to $\mathbf{\Gamma}(Y)$.
 The tree of finite (possibly empty) sequences of natural numbers is denoted $\omega^{<\omega}$.
@@
 Next, we set up an appropriate descriptive set-theoretic framework for the class of separable infinite-dimensional free spaces.
-We shall consider the Polish space $0,\infty)^{\omega\times\omega}$ endowed with the pointwise topology, and the set $\mathcal{R}\subset[0,\infty)^{\omega\times\omega}$ of metrics on $\omega$.
+We shall consider the Polish space $0,\infty)^{\omega\times\omega}$ endowed with the pointwise topology, and the set $\mathcal{R}\subseteq[0,\infty)^{\omega\times\omega}$ of metrics on $\omega$.
 This space was introduced by Vershik in [vershik:98] and features again in e.g.
@@
 This space can be used to code the class of complete separable infinite metric spaces.
-Given such a space $(M,\rho)$, fix a dense sequence of distinct points $(x_{i})\subset M$ and consider $d\in\mathcal{R}$ given by $d(i,j)=\rho(x_{i},x_{j})$, $i,j\in\omega$; then $(M,\rho)\equiv M_{d}$, where $M_{d}$ is the completion of $(\omega,d)$.
+Given such a space $(M,\rho)$, fix a dense sequence of distinct points $(x_{i})\subseteq M$ and consider $d\in\mathcal{R}$ given by $d(i,j)=\rho(x_{i},x_{j})$, $i,j\in\omega$; then $(M,\rho)\equiv M_{d}$, where $M_{d}$ is the completion of $(\omega,d)$.
 Conversely any such completion is separable and infinite.
@@
 As a change of base point does not alter the isometric structure of the corresponding free space, we shall always regard $0\in\omega$ as the base point.
-Bossard coded the class of separable infinite-dimensional Banach spaces as a Borel subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure.
-The authors of [cuth:doucha:kurka:22a, cuth:doucha:kurka:22b] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$.
+Bossard coded the class of separable infinite-dimensional Banach spaces as a $\mathbf{\Delta}^{1}_{1}$ subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure.
+Following a programme initiated in [godefroy:saint-raymond:18], the authors of [cuth:doucha:kurka:22a, cuth:doucha:kurka:22b] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$.
 First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ whose canonical extension to $c_{00}$ is a norm.
@@
 Every separable and infinite-dimensional Banach space $(X,\left\|\cdot\right\|)$ has representatives in $\mathcal{B}$.
-Given a dense linearly independent sequence $(x_{i})\subset X$, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$.
+Given a dense linearly independent sequence $(x_{i})\subseteq X$, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$.
 Then $X\equiv X_{\mu}$, where $X_{\mu}$ is the completion of $(V,\mu)$; conversely, every such completion is a separable and infinite-dimensional Banach space.
@@
 However we shall work with Vershik’s space $\mathcal{R}$ for the reasons given above.
-Analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [vershik:04]*Theorems 1 and 3.
+Also, analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [vershik:04]*Theorems 1 and 3.
 Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [fonf:wojtaszczyk:08], could be regarded as the generic separable free space.
 A immediate consequence of this and [fonf:wojtaszczyk:08]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual.
-We mention this because Polish spaces of metrics have been used elsewhere to study approximation properties in free spaces [talimdjioski:23, smith:talimdjioski:24, talimdjioski:25].
-3. The main results
+This observation complements Theorem 4.1.
+3. Three main results
 Our first result will allow us to apply to $\mathcal{R}$ one of the observations about $\mathcal{B}$ recorded above.
@@
 is $\mathbf{\Sigma}^{1}_{1}$, whence $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. ∎
-Next, while not strictly necessary for what is to come, we determine the descriptive complexity of the classes of countably infinite complete discrete metric spaces and complete separable p1u metric spaces.
+Next, we determine the descriptive complexity of the classes of countably infinite complete discrete metric spaces and complete separable p1u metric spaces.
 Proposition 3.3.
@@
 Then $(d,(x_{i}))\in A:=\bigcap_{k\in\omega}\bigcup_{m\in\omega}A_{k,m}$ if and only if $(x_{i})\in\omega^{\omega}$ is Cauchy in $(\omega,d)$, and $(x_{i})\in B:=\bigcup_{m\in\omega}B_{m}$ if and only if $(x_{i})$ is eventually constant.
-Therefore $\mathcal{R}\setminus\mathcal{R}_{\mathrm{dis}}$ is $\mathbf{\Sigma}^{1}_{1}$ because it is the image of $A\setminus(\mathcal{R}\times B)\subset\mathcal{R}\times\omega^{\omega}$ under the projection onto the first coordinate.
+Therefore $\mathcal{R}\setminus\mathcal{R}_{\mathrm{dis}}$ is $\mathbf{\Sigma}^{1}_{1}$ because it is the image of $A\setminus(\mathcal{R}\times B)\subseteq\mathcal{R}\times\omega^{\omega}$ under the projection onto the first coordinate.
 ∎
@@
 Proposition 3.4 can be proved more directly using Kirchheim’s characterisation of p1u spaces mentioned above.
-Later, in Corollary 3.9, we upgrade $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$ to $\mathbf{\Pi}^{1}_{1}$-complete.
+Later, in Proposition 3.13, we upgrade $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$ to $\mathbf{\Pi}^{1}_{1}$-complete.
 For this and the main results we will need the following lemma, which is a tool for generating continuous maps from power sets to $\mathcal{R}$.
 Lemma 3.5.
-Let $(M,d)$ be a metric space, $A\subset\omega$, ${\phi_{0}}:{A}\to{M}$ and ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ injections with pairwise disjoint ranges, and let ${\tau}:{\omega\setminus A}\to{\Gamma}$, where $\Gamma$ is a set.
+Let $(M,d)$ be a metric space, $A\subseteq\omega$, ${\phi_{0}}:{A}\to{M}$ and ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ injections with pairwise disjoint ranges, and let ${\tau}:{\omega\setminus A}\to{\Gamma}$, where $\Gamma$ is a set.
 Then the map ${\Psi}:{2^{\Gamma}}\to{M^{\omega}}$ given by
@@
 Proof.
-To check continuity of $\Psi$, let $E\in 2^{\Gamma}$ and let $F\subset\omega$ be finite.
-Define the finite sets $G=\tau(F\setminus A)\cap E$ and $H=\tau(F\setminus A)\setminus E$; then $G\subset E$, $H\cap E=\varnothing$ and $\Psi(E^{\prime})(n)=\Psi(E)(n)$ whenever $n\in F$, $E^{\prime}\in 2^{\Gamma}$, $G\subset E^{\prime}$ and $H\cap E^{\prime}=\varnothing$.
+To check continuity of $\Psi$, let $E\in 2^{\Gamma}$ and let $F\subseteq\omega$ be finite.
+Define the finite sets $G=\tau(F\setminus A)\cap E$ and $H=\tau(F\setminus A)\setminus E$; then $G\subseteq E$, $H\cap E=\varnothing$ and $\Psi(E^{\prime})(n)=\Psi(E)(n)$ whenever $n\in F$, $E^{\prime}\in 2^{\Gamma}$, $G\subseteq E^{\prime}$ and $H\cap E^{\prime}=\varnothing$.
 Therefore $\Psi$ is continuous.
@@
 We will be interested in metrics $\rho$ on $\omega^{<\omega}$ satisfying the following property.
-Given $\varepsilon>0$ and an infinite branch $B\subset\omega^{<\omega}$, there exists $n\in\omega$ such that
+Given $\varepsilon>0$ and an infinite branch $B\subseteq\omega^{<\omega}$, there exists $n\in\omega$ such that
 $\rho(s,t)<\varepsilon$ whenever $s,t\in B$ and $|s|,|t|\geqslant n$. (3.1)
+In the next result and a number of others, $\overline{{M}}$ shall denote the completion of the metric space $M$.
 Lemma 3.7.
-Let $(M,\theta)$ be a separable metric space, let $(x_{i})$ be a dense sequence in $M$ whose elements repeat infinitely often, and let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.1).
-Then the metric $d$ on $\omega^{<\omega}$ defined by
-$d(s,t)=\rho(s,t)+\theta(x_{|s|},x_{|t|}),$
-has the property that $(M,\theta)$ embeds isometrically into the completion $\overline{{(B,d)}}$ of $(B,d)$ whenever $B\subset\omega^{<\omega}$ is an infinite branch.
+Let $M=(M,\theta)$ be a separable metric space, let $(x_{i})_{i\in\omega}$ be a dense sequence in $M$ whose elements repeat infinitely often, let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.1) and define a second metric $\eta$ on $\omega^{<\omega}$ by
+$\eta(s,t)=\rho(s,t)+\theta(x_{|s|},x_{|t|}).$
+Then whenever $T\in\mathbf{IF}$ and $B\subseteq T$ is an infinite branch, there exists an isometric embedding ${\phi}:{M}\to{\overline{{(B,\eta)}}\subseteq\overline{{(T,\eta)}}}$.
+Moreover, if $M$ is complete then there is a 1-Lipschitz retraction ${r}:{\overline{{(T,\eta)}}}\to{\phi(M)}$.
 Proof.
-Let $B\subset\omega^{<\omega}$ be an infinite branch.
-Then there exists $q\in\omega^{\omega}$ such that $B=\left\{{q\mathord{\upharpoonright}_{n}}\;:\;\,{n\in\omega}\right\}$.
-We define an isometric embedding ${\phi}:{M}\to{\overline{{(B,d)}}}$ as follows.
+Let $q\in\omega^{\omega}$ such that $B=\left\{{q\mathord{\upharpoonright}_{n}}\;:\;\,{n\in\omega}\right\}$.
+We define an isometric embedding ${\phi}:{M}\to{\overline{{(B,\eta)}}}$ as follows.
 Given $x\in M$, because the elements of $(x_{i})$ repeat infinitely often, there exist $i_{0}<i_{1}<i_{2}<\ldots$ such that $\theta(x_{i_{k}},x)\to 0$.
-Given that $\rho$ satisfies (3.1), $(q\mathord{\upharpoonright}_{i_{k}})$ is $\rho$-Cauchy, hence $(q\mathord{\upharpoonright}_{i_{k}})$ is also $d$-Cauchy.
-Let $\phi(x)=\lim_{k}q\mathord{\upharpoonright}_{i_{k}}\in\overline{{(B,d)}}$.
-This is well-defined because if $j_{0}<j_{1}<j_{2}<\ldots$ also satisfy $\theta(x_{j_{k}},x)\to 0$ then $d(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})\to 0$, again by (3.1).
-Finally, we check that $\phi$ is an isometric embedding.
+Given that $\rho$ satisfies (3.1), $(q\mathord{\upharpoonright}_{i_{k}})$ is $\rho$-Cauchy, hence $(q\mathord{\upharpoonright}_{i_{k}})$ is also $\eta$-Cauchy.
+Let $\phi(x)=\lim_{k}q\mathord{\upharpoonright}_{i_{k}}\in\overline{{(B,\eta)}}$.
+This is well-defined because if $j_{0}<j_{1}<j_{2}<\ldots$ also satisfy $\theta(x_{j_{k}},x)\to 0$ then $\eta(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})\to 0$, again by (3.1).
+Next, we check that $\phi$ is an isometric embedding.
 With $x$ and the $i_{k}$ as above, let $y\in M$ and this time let $j_{0}<j_{1}<j_{2}<\ldots$ such that $\theta(y,x_{j_{k}})\to 0$.
 Then once more by (3.1),
-$d(\phi(x),\phi(y))=\lim_{k}d(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\lim_{k}\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})=\theta(x,y).\qed$
+$\eta(\phi(x),\phi(y))=\lim_{k}\eta(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\lim_{k}\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})=\theta(x,y).$
+Finally, assume $M$ is complete.
+Define ${\zeta}:{T}\to{M}$ by $\zeta(s)=x_{|s|}$.
+This map is $1$-Lipschitz because $\theta(\zeta(s),\zeta(t))=\theta(x_{|s|},x_{|t|})\leqslant\eta(s,t)$, $s,t\in T$.
+Let ${r}:{\overline{{(T,\eta)}}}\to{\overline{{(B,\eta)}}}$ be the $1$-Lipschitz extension of $\phi\circ\zeta$ to $\overline{{(T,\eta)}}$.
+By the continuity of $r$ and the completeness of $M$, we have $r(\overline{{(T,\eta)}})\subseteq\phi(M)$.
+Moreover $r(\phi(x))=\phi(x)$ whenever $x\in M$.
+Indeed, given $i_{0}<i_{1}<i_{2}<\ldots$ such that $\theta(x_{i_{k}},x)\to 0$, we have $\phi(x)=\lim_{k}q\mathord{\upharpoonright}_{i_{k}}$, so $r(\phi(x))=\lim_{k}\phi(\zeta(q\mathord{\upharpoonright}_{i_{k}}))=\lim_{k}\phi(x_{i_{k}})=\phi(x)$.
+∎
 If $(M,\theta)$ above is perfect then there is no need to insist that the elements of $(x_{i})$ repeat infinitely often.
@@
 Lemma 3.8.
-Let $d$ be the metric from Lemma 3.7, where $(M,\theta)$ is arbitrary and $\rho$ is as above. Then $(\omega^{<\omega},d)$ is discrete, and $(T,d)$ is in addition complete whenever $T\in\mathbf{WF}$.
+Let $\eta$ be the metric from Lemma 3.7, where $(M,\theta)$ is arbitrary and $\rho$ is as above.
+Then $(\omega^{<\omega},\eta)$ is discrete, and in addition $(T,\eta)$ is complete whenever $T\in\mathbf{WF}$.
 Proof.
 Fix $s\in\omega^{<\omega}$. Given $t\in\omega^{<\omega}\setminus\{s\}$, by (3.2) we have
-$d(s,t)\geqslant\rho(s,t)=2^{-|s\wedge t|-1}\geqslant 2^{-|s|-1}.$
-Hence $d(s,t)\geqslant 2^{-|s|-1}$ whenever $t\in\omega^{<\omega}\setminus\{s\}$, meaning that $d$ is discrete.
-Now let $T\in\mathbf{WF}$ and let $(s_{n})\subset T$ be $d$-Cauchy.
-We find a strictly increasing finite sequence $r_{0}\prec r_{1}\prec r_{2}\prec\ldots$ of elements of $\omega^{<\omega}$ and infinite subsets $L_{0}\supset L_{1}\supset L_{2}\supset\ldots$ of $\omega$ such that $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$.
+$\eta(s,t)\geqslant\rho(s,t)=2^{-|s\wedge t|-1}\geqslant 2^{-|s|-1}.$
+Hence $\eta(s,t)\geqslant 2^{-|s|-1}$ whenever $t\in\omega^{<\omega}\setminus\{s\}$, meaning that $\eta$ is discrete.
+Now let $T\in\mathbf{WF}$ and let $(s_{n})\subseteq T$ be $\eta$-Cauchy.
+We find a strictly increasing finite sequence $r_{0}\prec r_{1}\prec r_{2}\prec\ldots$ of elements of $\omega^{<\omega}$ and infinite subsets $L_{0}\supseteq L_{1}\supseteq L_{2}\supseteq\ldots$ of $\omega$ such that $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$.
 Begin by setting $r_{0}=\varnothing$ and $L_{0}=\omega$.
-Given $r_{k}$ and $L_{k}$, either there exists $m\in\omega$ and an infinite subset $L_{k+1}\subset L_{k}$ such that $r_{k}^{\frown}m\preccurlyeq s_{n}$ for all $n\in L_{k+1}$, or not.
+Given $r_{k}$ and $L_{k}$, either there exists $m\in\omega$ and an infinite subset $L_{k+1}\subseteq L_{k}$ such that $r_{k}^{\frown}m\preccurlyeq s_{n}$ for all $n\in L_{k+1}$, or not.
 If the first case holds, set $r_{k+1}=r_{k}^{\frown}m$ and continue the process, else terminate it.
@@
 If $n_{1}\in H_{m_{1}}$ and $n_{2}\in H_{m_{2}}$ for distinct $m_{1},m_{2}\in\omega$, then $s_{n_{1}}\perp s_{n_{2}}$ and $s_{n_{1}}\wedge s_{n_{2}}=r_{k}$, giving
-$d(s_{n_{1}},s_{n_{2}})\geqslant\rho(s_{n_{1}},s_{n_{2}})=2^{-|r_{k}|-1}.$
-Because $(s_{n})$ is $d$-Cauchy, it follows that $H_{m}$ is empty for all but finitely many $m\in\omega$, and thus $s_{n}=r_{k}$ for infinitely many $n\in L_{k}$.
-Again because $(s_{n})$ is $d$-Cauchy, we conclude that it must converge to $r_{k}$.
-This proves that $(T,d)$ is complete.
+$\eta(s_{n_{1}},s_{n_{2}})\geqslant\rho(s_{n_{1}},s_{n_{2}})=2^{-|r_{k}|-1}.$
+Because $(s_{n})$ is $\eta$-Cauchy, it follows that $H_{m}$ is empty for all but finitely many $m\in\omega$, and thus $s_{n}=r_{k}$ for infinitely many $n\in L_{k}$.
+Again because $(s_{n})$ is $\eta$-Cauchy, we conclude that it must converge to $r_{k}$.
+This proves that $(T,\eta)$ is complete.
 ∎
-The proofs of Theorem 1.1 and Theorem 1.2 use Bossard’s reduction argument [bossard:02]*Theorem 3.2.
+We make use of the last three lemmas in the next result. Lemmas 3.9 and 3.10 below are our main tools.
+Lemma 3.9.
+Let $M=(M,\theta)$ be a separable metric space.
+There exists a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ whenever $T\in\mathbf{WF}$ and $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$ whenever $T\in\mathbf{IF}$.
+Proof.
+Let $(\omega^{<\omega},\eta)$ be the metric space obtained from Lemma 3.7 applied to $M$.
+Let $W=\omega^{<\omega}\cup(\omega^{<\omega}\cup\{0\})$.
+We extend $\eta$ to a metric on $W$, again labelled $\eta$, by setting
+$\eta(x,y)=\begin{cases}0&\text{if }x,y\notin\omega^{<\omega},\,x=y,\\ 1&\text{if }x,y\notin\omega^{<\omega},\,x\neq y,\\ 1+\eta(\varnothing,x)&\text{if }x\in\omega^{<\omega},\,y\notin\omega^{<\omega},\\ 1+\eta(\varnothing,y)&\text{if }x\notin\omega^{<\omega},\,y\in\omega^{<\omega}.\end{cases}$
+It is routine to verify that this extension satisfies the triangle inequality.
+Set $A=\phi_{0}=\varnothing$, $\Gamma=\omega^{<\omega}$, let ${\tau}:{\omega}\to{\Gamma}$ be a bijection, and define injections ${\phi_{1},\phi_{2}}:{\omega}\to{W}$ by $\phi_{1}(n)=(\tau(n),0)$ and $\phi_{2}(n)=\tau(n)$, $n\in\omega$.
+Let ${\Delta}:{2^{\Gamma}}\to{\mathcal{R}}$ be the continuous map furnished by Lemma 3.5, which we then restrict to $\mathbf{Tr}$.
+We observe that $\Psi(T)(\omega)=T\cup((\omega^{<\omega}\setminus T)\times\{0\})$ for all $T\in\mathbf{Tr}$.
+Let $T\in\mathbf{WF}$.
+Using Lemma 3.8, $(T,\eta)$ is discrete and complete.
+Hence $(\omega,\Delta(T))$, which is isometric to $(\Psi(T)(\omega),\eta)$, is also discrete and complete because every point in $(\omega^{<\omega}\setminus T)\times\{0\}$ is $\eta$-separated from everything else by at least $1$.
+Therefore $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$.
+Now let $T\in\mathbf{IF}$. We claim that $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$. First of all, the function ${r^{\prime}}:{\Psi(T)(\omega)}\to{T}$, defined by
+$r^{\prime}(x)=\begin{cases}x&\text{if }x\in T,\\ \varnothing&\text{if }x\notin T,\end{cases}$
+is a $1$-Lipschitz retraction of $(\Psi(T)(\omega),\eta)$ onto $(T,\eta)$.
+We extend this to a $1$-Lipschitz retraction ${r^{\prime}}:{\overline{{(\Psi(T)(\omega),\eta)}}}\to{\overline{{(T,\eta)}}}$.
+According to Lemma 3.7, there is an isometric embedding ${\phi}:{M}\to{\overline{{(T,\eta)}}}$ and a $1$-Lipschitz retraction ${r}:{\overline{{(T,\eta)}}}\to{\phi(M)}$.
+Therefore ${r\circ r^{\prime}}:{\overline{{(\Psi(T)(\omega),\eta)}}}\to{\phi(M)}$ is a 1-Lipschitz retraction.
+Since $M_{\Delta(T)}\equiv\overline{{(\Psi(T)(\omega),\eta)}}$, our claim is proved.
+∎
+The final lemma below uses Bossard’s reduction argument [bossard:02]*Theorem 3.2; Theorems 1.11.3 emerge as relatively straightforward consequences of it.
+Lemma 3.10.
+Let $\mathcal{A}\subseteq\mathcal{R}$ have the property that $d^{\prime}\in\mathcal{A}$ whenever $d^{\prime}\in\mathcal{R}$ and $M_{d^{\prime}}$ is isometric to a $1$-Lipschitz retract of $M_{d}$ for some $d\in\mathcal{A}$.
+Assume that $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$ is $\mathbf{\Sigma}^{1}_{1}$.
+Then $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}=\mathcal{R}$.
+Proof.
+Let $d\in\mathcal{R}$.
+Apply Lemma 3.9 to $M:=M_{d}$.
+The assumption $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$ is $\mathbf{\Sigma}^{1}_{1}$ implies $\Delta^{-1}(\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A})$ is also $\mathbf{\Sigma}^{1}_{1}$.
+As $\mathbf{WF}\subseteq\Delta^{-1}(\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A})$ and $\mathbf{WF}$ is $\mathbf{\Pi}^{1}_{1}$-complete, by Suslin’s theorem there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$.
+Because $T\in\mathbf{IF}$, $M_{d}$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$.
+Given our assumption about $\mathcal{A}$ and the fact that $\mathcal{R}_{\mathrm{dis}}$ also enjoys this property, it follows that $d\in\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$.
+Therefore $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}=\mathcal{R}$.
+∎
 Proof of Theorem 1.1.
@@
 By Proposition 3.2, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$.
-Assume that $\mathcal{R}_{\mathrm{dis}}\subset\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
-Let $\theta$ be the usual metric on $B_{c_{0}}$ inherited from $c_{0}$, and let $(\omega^{<\omega},d)$ be the metric space obtained from Lemma 3.7 applied to $(B_{c_{0}},\theta)$.
-It is evident that $d\leqslant\frac{5}{2}$.
-Let $M=\omega^{<\omega}\cup(\omega^{<\omega}\times\{0\})$ and extend $d$ to a metric on $M$ by setting $d(x,y)=\frac{5}{2}$ whenever $x\neq y$ and $\{x,y\}\setminus\omega^{<\omega}$ is non-empty (it is straightforward to check that the extended function satisfies the triangle inequality).
-Set $A=\phi_{0}=\varnothing$, $\Gamma=\omega^{<\omega}$, let ${\tau}:{\omega}\to{\Gamma}$ be a bijection, and define injections ${\phi_{1},\phi_{2}}:{\omega}\to{M}$ by $\phi_{1}(n)=(\tau(n),0)$ and $\phi_{2}(n)=\tau(n)$, $n\in\omega$.
-Let ${\Delta}:{2^{\Gamma}}\to{\mathcal{R}}$ be the continuous map furnished by Lemma 3.5, which we then restrict to $\mathbf{Tr}$.
-Let $T\in\mathbf{Tr}$.
-We observe that $\Psi(T)(\omega)=T\cup((\omega^{<\omega}\setminus T)\times\{0\})$.
-Using Lemma 3.8, $(T,d)$ is discrete and complete if $T\in\mathbf{WF}$.
-Hence $(\omega,\Delta(T))$, which is isometric to $(\Psi(T)(\omega),d)$, is also discrete and complete because every point in $(\omega^{<\omega}\setminus T)\times\{0\}$ is $d$-separated from everything else by $\frac{5}{2}$.
-Therefore $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$.
-Now let $T\in\mathbf{IF}$ and let $B\subset T$ be an infinite branch.
-Then the completion $M_{\Delta(T)}$ of $(\omega,\Delta(T))$ contains an isometric copy of the completion of $(B,d)$.
-By Lemma 3.7, $B_{c_{0}}\hookrightarrow M_{\Delta(T)}$.
-As $\mathcal{R}_{\mathrm{dis}}\subset\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$, we have $\mathbf{WF}=\Delta^{-1}(\mathcal{R}_{\mathrm{dis}})\subset\Delta^{-1}(\langle X\rangle^{\mathcal{R}}_{\hookrightarrow})$.
-Since $\mathbf{WF}$ is $\mathbf{\Pi}^{1}_{1}$-complete and $\Delta^{-1}(\langle X\rangle^{\mathcal{R}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$, there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
-Therefore $\mathcal{F}({B_{c_{0}}})\hookrightarrow\mathcal{F}({M_{d}})\hookrightarrow X$.
-Because $Y\hookrightarrow\mathcal{F}({B_{c_{0}}})$ whenever $Y$ is a separable Banach space ([dutrieux:ferenczi:06]*p.
-1042 and [kaufmann:15]*Corollary 3.3), the conclusion follows.
+Let $d^{\prime}\in\mathcal{R}$ and $d\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$.
+By [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})\hookrightarrow\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
+Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
+By Lemma 3.10, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$.
+As $Y\hookrightarrow\mathcal{F}({c_{0}})$ whenever $Y$ is a separable Banach space ([dutrieux:ferenczi:06]*p.
+1042) and $c_{0}\equiv M_{d}$ for some $d\in\mathcal{R}$, it follows that $Y\hookrightarrow X$ for all such $Y$.
 ∎
@@
 By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$.
-By repeating part of the proof above, this time applying Lemma 3.7 to $\mathbb{U}$, we obtain a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ if $T\in\mathbf{WF}$ and $\mathbb{U}\hookrightarrow M_{\Delta(T)}$ otherwise.
-By the last part of the proof above, there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$, which implies $\mathbb{U}\hookrightarrow M_{\Delta(T)}\hookrightarrow M$.
+Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
+By Lemma 3.10, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$.
+By considering e.g.
+$\mathbb{U}$ or $c_{0}$ (the latter being Lipschitz universal by Aharoni’s theorem, see e.g.
+[benyamini:lindenstrauss:00]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal.
 ∎
-Corollary 3.9.
+Proof of Theorem 1.3.
+Using [ghawadrah:17], [cuth:doucha:kurka:22a]*Theorem 8 and Proposition 3.1, the set
+[…]
+is $\mathbf{\Delta}^{1}_{1}$.
+Let $d^{\prime}\in\mathcal{R}$ and $d\in\mathcal{R}_{\mathrm{BAP}}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$.
+Again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\mathcal{R}_{\mathrm{BAP}}$.
+For a contradiction assume $\mathcal{R}_{\mathrm{dis}}\subseteq\mathcal{R}_{\mathrm{BAP}}$.
+Then $\mathcal{R}_{\mathrm{BAP}}=\mathcal{R}$ by Lemma 3.10.
+However, this is false because we can pick $d\in\mathcal{R}$ such that $M_{d}$ is isometric to a separable Banach space that fails the AP; as is pointed out in the Introduction, $d\notin\mathcal{R}_{\mathrm{BAP}}$ by [godefroy:kalton:03]*Theorem 3.1.
+Now pick $d\in\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$.
+Again, as stated in the Introduction, $\mathcal{F}({M_{d}})$ has the AP by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8.
+That $\mathcal{F}({M_{d}})$ cannot be isomorphic to a dual space is a simple consequence of the classical result of Grothendieck which asserts that a separable Banach space has the MAP whenever it has the AP and is isometric to a dual space.
+∎
+Remark 3.11.
+The set $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ is $\mathbf{\Pi}^{1}_{1}$ by Proposition 3.3 and the fact that $\mathcal{R}_{\mathrm{BAP}}$ is $\mathbf{\Delta}^{1}_{1}$.
+Theorem 1.3 shows that $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ is non-empty.
+By a simple adjustment to the argument we see moreover that $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ cannot be $\mathbf{\Delta}^{1}_{1}$: if it was $\mathbf{\Delta}^{1}_{1}$ then $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{R}_{\mathrm{BAP}}$ would be as well, thus it would be equal to $\mathcal{R}$ by Lemma 3.10, which again is false.
+We make some remarks based on this observation.
+  1. (1)
+    Given metric spaces $M$ and $N$, let us write $M\stackrel{{\scriptstyle u}}{{\hookrightarrow}}N$ if $M$ is uniformly homeomorphic to a subspace of $N$. Similarly to [clemens:12]*Lemma 4, the set
+    $\{(d,d^{\prime})\in\mathcal{R}\times\mathcal{R}\,:\,M_{d}\stackrel{{\scriptstyle u}}{{\hookrightarrow}}M_{d^{\prime}}\}$
+    is $\mathbf{\Sigma}^{1}_{1}$ and thus so is $\langle M\rangle^{\mathcal{R},\,u}_{\hookrightarrow}:=\{d\in\mathcal{R}\,:\,M_{d}\stackrel{{\scriptstyle u}}{{\hookrightarrow}}M\}$ whenever $M$ is complete and separable.
+Evidently $\langle M_{d}\rangle^{\mathcal{R},\,u}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{dis}}$ whenever $d\in\mathcal{R}_{\mathrm{dis}}$.
+Hence, given a countable subset $C\subseteq\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$, the set $\Gamma:=\bigcup_{d\in C}\langle M_{d}\rangle^{\mathcal{R},\,u}_{\hookrightarrow}\setminus\mathcal{R}_{\mathrm{BAP}}$ is $\mathbf{\Sigma}^{1}_{1}$ and is included in $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$, which is $\mathbf{\Pi}^{1}_{1}$ but not $\mathbf{\Delta}^{1}_{1}$ from above.
+By Suslin’s theorem there exists $d^{\prime}\in\mathcal{R}_{\mathrm{dis}}\setminus(\mathcal{R}_{\mathrm{BAP}}\cup\Gamma)$.
+We can apply transfinite induction to this fact to obtain a sequence $(d_{\alpha})_{\alpha<{\omega_{1}}}$ of elements of $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ such that $M_{d_{\beta}}\not\stackrel{{\scriptstyle u}}{{\hookrightarrow}}M_{d_{\alpha}}$ whenever $\alpha<\beta<\omega_{1}$.
+I don’t know if the $d_{\alpha}$ can be constructed so that $\mathcal{F}({M_{d_{\alpha}}})$ is not isomorphic to $\mathcal{F}({M_{d_{\beta}}})$ whenever $\alpha\neq\beta$.
+  2. (2)
+    We can replace $\mathcal{R}_{\mathrm{dis}}$ by $\mathcal{R}_{\mathrm{p1u}}$ to show that the set $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ (which is $\mathbf{\Pi}^{1}_{1}$ by Proposition 3.4) is not $\mathbf{\Delta}^{1}_{1}$.
+No changes to the proofs above are needed.
+Given $d\in\mathcal{R}_{\mathrm{p1u}}$, we have $\langle\mathcal{F}({M_{d}})\rangle^{\mathcal{R}}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{p1u}}$ by [aliaga:gartand:petitjean:prochazka:22]*Theorem 4.6.
+Therefore, by arguing as above, there is a sequence $(d_{\alpha})_{\alpha<\omega_{1}}$ of elements of $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ such that $\mathcal{F}({M_{d_{\beta}}})\not\hookrightarrow\mathcal{F}({M_{d_{\alpha}}})$ whenever $\alpha<\beta<\omega_{1}$.
+I don’t know if the spaces $\mathcal{F}({M_{d_{\alpha}}})$ have the AP.
+  3. (3)
+    While the above results show that there are many examples of complete separable p1u spaces whose free spaces fail the BAP, they don’t provide an explicit description of any of them.
+It could be instructive to obtain concrete examples.
+I am grateful to G. Lancien for suggesting the following problem which Theorem 1.3 leaves open.
+Problem 3.12.
+Let $M$ be a complete separable p1u metric space. Does there exist a separable Banach space $X$ such that $\mathcal{F}({M})\hookrightarrow X^{*}$?
+We finish this section with two direct applications of Lemma 3.9.
+Proposition 3.13.
 The sets $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$ are $\mathbf{\Pi}^{1}_{1}$-complete.
 Proof.
-The maps $\Delta$ in the proofs of Theorems 1.1 and 1.2 witness the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$. ∎
-We finish by making a few remarks about Theorems 1.1 and 1.2.
-First, it is clear that $M_{\Delta(T)}$ in the proof of Theorem 1.1 is always bounded, so if we require that $X$ is universal for the class of free spaces over countable complete discrete spaces that are in addition bounded, the outcome is the same.
-Of course, in the proof of Theorem 1.2 we can replace $\mathbb{U}$ by any separable Lipschitz universal space such as $c_{0}$ (by Aharoni’s theorem, see e.g.
-[benyamini:lindenstrauss:00]*Theorem 7.11).
-If we add boundedness to the list of properties in Theorem 1.2 then $M$ is Lipschitz universal for all bounded separable metric spaces.
-This is because we can replace $\mathbb{U}$ in the proof e.g.
-by any ball in $\mathbb{U}$ (see [vershik:98]*p.
-922) or any ball in $c_{0}$.
-Finally, we can add the property of being uniformly concave to the lists above and obtain the same conclusions.
-According to [weaver:18]*Definition 3.33, $(M,d)$ is uniformly concave if, given distinct $x,z\in M$ and $\varepsilon>0$, there exists $\delta>0$ such that $d(x,z)\leqslant d(x,y)+d(y,z)-\delta$ whenever $d(x,y),d(y,z)\geqslant\varepsilon$.
-Uniformly concave metric spaces are relevant to the extremal structure of free spaces (see e.g.
-[weaver:18]*Theorem 3.39).
-It can be shown that $(\omega,\Delta(T))=M_{\Delta(T)}$ is uniformly concave whenever $T\in\mathbf{WF}$.
-This holds because $(\omega^{<\omega},\rho)$ is uniformly concave (with $\delta$ independent of $\varepsilon$).
-To see this we require a quick lemma.
-Lemma 3.10.
-If $s,t,u\in\omega^{<\omega}$ and $s\wedge u\prec s\wedge t$ then $s\wedge u=t\wedge u$.
+Recall that, by Propositions 3.3 and 3.4, these sets are $\mathbf{\Pi}^{1}_{1}$.
+Select any complete separable non-p1u metric space $M$.
+Then the map $\Delta$ furnished by Lemma 3.9 applied to $M$ witnesses the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$.
+∎
+In the final result of this section, we obtain a lower bound on the complexity of the set of metrics in $\mathcal{R}$ whose free spaces satisfy the AP.
+The complexity calculation is completed in Section 4.
+Proposition 3.14.
+[…]
 Proof.
-As $s\wedge u\preccurlyeq u$ and $s\wedge u\prec s\wedge t\preccurlyeq t$, we have $s\wedge u\preccurlyeq t\wedge u$.
-Because $s\wedge t,t\wedge u\preccurlyeq t$, they are comparable.
-However, if $s\wedge t\preccurlyeq t\wedge u\preccurlyeq u$ then $s\wedge t\preccurlyeq s\wedge u$, which is not the case.
-Thus $t\wedge u\prec s\wedge t\preccurlyeq s$, giving $t\wedge u\preccurlyeq s\wedge u$.
+Apply Lemma 3.9 to a separable Banach space $M$ that fails the AP.
+If $T\in\mathbf{WF}$ then $\mathcal{F}({M_{\Delta(T)}})$ has the AP again by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8.
+If $T\in\mathbf{IF}$ then $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$.
+Once again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{\Delta(T)}})$.
+Since $\mathcal{F}({M})$ fails the AP, so must $\mathcal{F}({M_{\Delta(T)}})$.
+Therefore $\Delta$ witnesses the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{AP}}$.
 ∎
-Now let $s,t,u\in\omega^{<\omega}$ be distinct and set $\delta=\min\{2^{-|s|-1},2^{-|u|-1}\}>0$.
-We will show that $\rho(s,u)\leqslant\rho(s,t)+\rho(t,u)-\delta$, thus demonstrating that $(\omega^{<\omega},\rho)$ is uniformly concave.
-If $|s\wedge t|\leqslant|s\wedge u|$ then
-$\rho(s,t)+\rho(t,u)=2^{-|s\wedge t|-1}+2^{-|t\wedge u|-1}\geqslant 2^{-|s\wedge u|-1}+2^{-|u|-1}\geqslant\rho(s,u)+\delta;$
-and we get the same result if $|t\wedge u|\leqslant|s\wedge u|$.
-On the other hand, if $|s\wedge u|<|s\wedge t|$ then $s\wedge u\prec s\wedge t$ because these elements are comparable, thus $s\wedge u=t\wedge u$ by Lemma 3.10.
-Then we estimate
-$\rho(s,t)+\rho(t,u)=2^{-|s\wedge t|-1}+2^{-|t\wedge u|-1}\geqslant 2^{-|s|-1}+2^{|s\wedge u|-1}\geqslant\delta+\rho(s,u).$
-We reach the same conclusion if $|s\wedge u|<|t\wedge u|$.
+4. Complexity and the approximation property
+This section is solely concerned with determining the complexity of two classes of Banach spaces having the AP.
+The proof of the result below is an adaptation of an argument from [ghawadrah:17].
+As above, we frame the result in terms of the Polish space $\mathcal{B}$ instead of the Effros-Borel space $\mathcal{SB}$.
+Theorem 4.1.
+[…]
+Proof.
+Having in mind Propositions 3.1 and 3.14, it is enough to prove that $\mathcal{B}_{\mathrm{AP}}$ is $\mathbf{\Pi}^{1}_{1}$.
+Let $(e_{i})_{i\in\omega}\subseteq V$ denote the standard basis of $c_{00}$; then $(e_{i})_{i\in\omega}$ is a linearly independent sequence having dense span in $X_{\mu}$ whenever $\mu\in\mathcal{B}$.
+Given $R\in\omega$, let $V_{R}=\mathbb{Q}$-$\operatorname{span}(e_{i})_{i<R}\subseteq V$.
+Define the closed subset
+$RC=\left\{{(\mu,x)\in\mathcal{B}\times V^{\omega\times\omega}}\;:\;\,{\mu(x(k,p)-x(k,q))\leqslant 2^{-p}\text{ for all }k,p,q\in\omega,\,q\geqslant p}\right\}$
+of the Polish space $\mathcal{B}\times V^{\omega\times\omega}$ (where $V$ has the discrete topology). Then, given finite $F\subseteq\omega$ and $n,k,j\in\omega$, define the closed subset
+$TB_{n,F,k,j}=\left\{{(\mu,x)\in\mathcal{B}\times V^{\omega\times\omega}}\;:\;\,{\mu(x(k,q)-x(j,q))\leqslant 2^{-n}\text{ for all }q,n\in\omega,\,q\geqslant n+1}\right\}$
+of $\mathcal{B}\times V^{\omega\times\omega}$ and the $\mathbf{\Pi}^{0}_{3}$ subset
+$TB=\bigcap_{n\in\omega}\bigcup_{\begin{subarray}{c}F\subseteq\omega\\ \mathrm{finite}\end{subarray}}\bigcap_{k\in\omega}\bigcup_{j\in F}TB_{n,F,k,j}.$
+Next, given $n,L,R,N\in\omega$, $v=(v_{i})_{i<N}\in V_{R}^{N}$ and $\alpha=(\alpha_{i})_{i<N}\in\mathbb{Q}^{N}$, define the closed subset
+$FR_{L,N,v,\alpha}=\left\{{\mu\in\mathcal{B}}\;:\;\,{\mu\bigg(\sum_{i<N}\alpha_{i}v_{i}\bigg)\leqslant L\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg)}\right\}$
+of $\mathcal{B}$, and the closed subset
+$AP_{n,L,N,v}=$ $\bigg\{(\mu,x)\in\mathcal{B}\times V^{\omega\times\omega}\,:\,q\leqslant N\text{ and }x(k,q)\in V_{N}\text{ implies}$
+$\mu\bigg(x(k,q)-\sum_{i<N}x(k,q)(i)v_{i}\bigg)\leqslant 2^{-n}+(L+2)\cdot 2^{-q}\bigg\}$
+of $\mathcal{B}\times V^{\omega\times\omega}$. Then define the $\mathbf{\Pi}^{0}_{5}$ subset
+$AP=\bigcap_{n\in\omega}\bigcup_{L,R\in\omega}\bigcap_{N\in\omega}\bigcup_{v\in V_{R}^{N}}\bigg(AP_{n,L,N,v}\cap\bigcap_{\alpha\in\mathbb{Q}^{N}}(FR_{L,N,v,\alpha}\times V^{\omega\times\omega})\bigg).$
+of $\mathcal{B}\times V^{\omega\times\omega}$.
+Now let $\mathcal{A}\subseteq\mathcal{B}$ be the $\mathbf{\Sigma}^{1}_{1}$ image of the projection onto the first coordinate of $(RC\cap TB)\setminus AP\subseteq\mathcal{B}\times V^{\omega\times\omega}$.
+We show that $\mathcal{B}_{\mathrm{AP}}=\mathcal{B}\setminus\mathcal{A}$ and thus $\mathcal{B}_{\mathrm{AP}}$ is $\mathbf{\Pi}^{1}_{1}$.
+Let $\mu\in\mathcal{B}\setminus\mathcal{A}$.
+We prove that $X_{\mu}$ has the AP, i.e.
+$\mu\in\mathcal{B}_{\mathrm{AP}}$.
+Let $K\subseteq X_{\mu}$ be compact and let $\varepsilon>0$.
+We can assume that $\mu(z)\leqslant 1$ whenever $z\in K$.
+Let $(z_{k})_{k\in\omega}$ be a dense sequence in $K$.
+Then pick $x\in V^{\omega\times\omega}$ such that $\mu(x(k,q)-z_{k})\leqslant 2^{-q-1}$ for all $k,q\in\omega$.
+We verify that $(\mu,x)\in RC\cap TB$.
+Evidently $(\mu,x)\in RC$.
+Now let $n\in\omega$.
+Because $(z_{k})_{k\in\omega}$ is totally bounded, there is finite $F\subseteq\omega$ such that whenever $k\in\omega$, there exists $j\in F$ satisfying $\mu(z_{k}-z_{j})\leqslant 2^{-n-1}$.
+By the choice of $x$,
+$\mu(x(k,q)-x(j,q))\leqslant 2^{-q}+2^{-n-1}\leqslant 2^{-n},\quad q\geqslant n+1.$
+It follows that $(\mu,x)\in TB$.
+As $\mu\notin\mathcal{A}$ and $(\mu,x)\in RC\cap TB$, it must be that $(\mu,x)\in AP$.
+Fix $n\in\omega$ such that $2^{-n}\leqslant\varepsilon$.
+Given $N\in\omega$, by equivalence of norms on finite-dimensional spaces, there exists $\theta_{N}>0$ such that
+$\sum_{i<N}|\alpha_{i}|\leqslant\theta_{N}\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg),\quad(\alpha_{i})_{i<N}\in\mathbb{R}^{N}.$ (4.1)
+As $(\mu,x)\in AP$, there exist $L,R\in\omega$ and $v^{(N)}\in V_{R}^{N}$, $N\in\omega$, such that
+$(\mu,x)\in AP_{n,L,N,v^{(N)}}\cap\bigcap_{\alpha\in\mathbb{Q}^{N}}(FR_{L,N,v^{(N)},\alpha}\times V^{\omega\times\omega}),\quad N\in\omega.$
+Given $\alpha\in\mathbb{Q}^{N}$, as $\mu\in FR_{L,N,v^{(N)},\alpha}$,
+$\mu\bigg(\sum_{i<N}\alpha_{i}v_{i}^{(N)}\bigg)\leqslant L\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg).$ (4.2)
+In particular, $\mu(v^{(N)}_{i})\leqslant L\mu(e_{i})$ whenever $i<N$.
+Since each $(v_{i}^{(N)})$, $i\in\omega$, is a bounded sequence in $V_{R}\subseteq\operatorname{span}(e_{j})_{j<R}$, by a diagonalisation argument there are $w_{i}\in\operatorname{span}(e_{j})_{j<R}$ and $N_{0}<N_{1}<N_{2}<\ldots$ such that $\mu\big(v^{(N_{\ell})}_{i}-w_{i}\big)\to 0$ for all $i\in\omega$.
+Using (4.2) and the continuity of $\mu$, given $\alpha\in\mathbb{Q}^{N}$,
+$\mu\bigg(\sum_{i<N}\alpha_{i}w_{i}\bigg)\leqslant L\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg).$
+Hence we can define a bounded $\mathbb{Q}$-linear operator ${T}:{V}\to{\operatorname{span}(e_{i})_{i<R}}$ by setting $Te_{i}=w_{i}$, $i\in\omega$, and then extending to finite $\mathbb{Q}$-linear combinations of the $e_{i}$.
+By equicontinuity we extend it again to a bounded linear operator on $X_{\mu}$ satisfying $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$ and $\mu(T)\leqslant L$.
+By taking a further subsequence of natural numbers, also labelled $(N_{\ell})$, we can ensure that
+$\mu\big(v_{i}^{(N_{\ell})}-w_{i}\big)\leqslant\frac{2^{-\ell-1}}{\theta_{\ell}},\quad i<\ell.$ (4.3)
+Now let $k,q\in\omega$.
+Pick $\ell\in\omega$ large enough so that $q\leqslant\ell$ and $x(k,q)\in V_{\ell}$.
+Then $\ell\leqslant N_{\ell}$ implies $q\leqslant N_{\ell}$ and $x(k,q)\in V_{N_{\ell}}$.
+Because $(\mu,x)\in AP_{n,L,N_{\ell},v^{(N_{\ell})}}$,
+$\mu\bigg(x(k,q)-\sum_{i<N_{\ell}}x(k,q)(i)v_{i}^{(N_{\ell})}\bigg)\leqslant 2^{-n}+(L+2)\cdot 2^{-q}.$ (4.4)
+We estimate
+$\mu\bigg(\sum_{i<\ell}x(k,q)(i)\big(v_{i}^{(N_{\ell})}-w_{i}\big)\bigg)$ $\leqslant\sum_{i<\ell}|x(k,q)(i)|\mu\big(v_{i}^{(N_{\ell})}-w_{i}\big)$
+$\leqslant\frac{2^{-\ell-1}}{\theta_{\ell}}\sum_{i<\ell}|x(k,q)(i)|$ by (4.3)
+$\leqslant 2^{-\ell-1}\mu\bigg(\sum_{i<\ell}x(k,q)(i)e_{i}\bigg)$ by (4.1)
+$=2^{-\ell-1}\mu(x(k,q))$ as $x(k,q)\in V_{\ell}$
+$\leqslant 2^{-\ell}\leqslant 2^{-q}.$
+The penultimate inequality above holds because $\mu(x(k,q))\leqslant\mu(z_{k})+2^{-1}\leqslant\frac{3}{2}$ for all $k,q\in\omega$. Therefore
+$\mu(x(k,q)-Tx(k,q))$ $=\mu\bigg(x(k,q)-\sum_{i<\ell}x(k,q)(i)w_{i}\bigg)$ as $x(k,q)\in V_{\ell}$
+$\leqslant\mu\bigg(x(k,q)-\sum_{i<\ell}x(k,q)(i)v_{i}^{(N_{\ell})}\bigg)+2^{-q}$
+$=\mu\bigg(x(k,q)-\sum_{i<N_{\ell}}x(k,q)(i)v_{i}^{(N_{\ell})}\bigg)+2^{-q}$ as $\ell\leqslant N_{\ell}$
+$\leqslant 2^{-n}+(L+3)\cdot 2^{-q}$ by (4.4).
+Fixing $k\in\omega$ and letting $q\to\infty$ yields $\mu(z_{k}-Tz_{k})\leqslant 2^{-n}$.
+Since this holds for all $k\in\omega$, by continuity $\mu(z-Tz)\leqslant 2^{-n}\leqslant\varepsilon$ for all $z\in K$.
+Therefore $X_{\mu}$ has the AP.
+Conversely, assume that $\mu\in\mathcal{B}_{\mathrm{AP}}$.
+Let $(\mu,x)\in RC\cap TB$.
+We show that $(\mu,x)\in AP$ and hence $\mu\in\mathcal{B}\setminus\mathcal{A}$.
+Let $n\in\omega$.
+Given $k\in\omega$, let $z_{k}=\lim_{q}x(k,q)\in X_{\mu}$.
+As $(\mu,x)\in TB$ as well, it is easy to see that $\left\{{z_{k}}\;:\;\,{k\in\omega}\right\}$ is totally bounded and thus
+$K:=\overline{{\left\{{z_{k}}\;:\;\,{k\in\omega}\right\}}}\subseteq X_{\mu}$
+is compact.
+As $X_{\mu}$ has the AP, there exists a bounded linear finite-rank operator ${T}:{X_{\mu}}\to{X_{\mu}}$ such that $\mu(z-Tz)<2^{-n}$ for all $z\in K$.
+Fix $L\in\omega$ such that $\mu(T)<L$.
+By a perturbation argument, we can assume there exists $R\in\omega$ such that $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$.
+Let $B>0$ such that $\mu(z)\leqslant B$ for all $z\in K$.
+Given $N\in\omega$, pick $\theta_{N}>0$ to satisfy (4.1) above and choose $v\in V^{N}_{R}$ such that
+$\mu(v_{i}-Te_{i})\leqslant\frac{1}{\theta_{N}}\min\bigg\{\frac{2^{-N}}{B+\frac{1}{2}},L-\mu(T)\bigg\},\quad i<N.$ (4.5)
+Then, given $\alpha\in\mathbb{Q}^{N}$,
+$\mu\bigg(\sum_{i<N}\alpha_{i}v_{i}\bigg)$ $\leqslant\mu\bigg(\sum_{i<N}\alpha_{i}Te_{i}\bigg)+\sum_{i<N}|\alpha_{i}|\mu(v_{i}-Te_{i})$
+$\leqslant\mu(T)\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg)+(L-\mu(T))\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg)$ by (4.1), (4.5)
+$=L\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg).$
+We conclude that
+$(\mu,x)\in\bigcap_{\alpha\in\mathbb{Q}^{N}}(FR_{L,N,v,\alpha}\times V^{\omega\times\omega}).$
+We also need to show that $(\mu,x)\in AP_{n,L,N,v}$. Assume that $q\leqslant N$ and $x(k,q)\in V_{N}$. As $(\mu,x)\in RC$ and $\mu(T)\leqslant L$, we have
+$\mu(x(k,q)-z_{k})\leqslant 2^{-q}\quad\text{and}\quad\mu(Tx(k,q)-Tz_{k})\leqslant L\cdot 2^{-q}.$ (4.6)
+Moreover,
+$\mu\bigg(Tx(k,q)-\sum_{i<N}x(k,q)(i)v_{i}\bigg)$ $\leqslant\sum_{i<N}|x(k,q)(i)|\mu(Te_{i}-v_{i})$ as $x(k,q)\in V_{N}$
+$\leqslant\frac{2^{-N}}{B+\frac{1}{2}}\mu\bigg(\sum_{i<N}x(k,q)(i)e_{i}\bigg)$ by (4.1), (4.5)
+$=\frac{2^{-N}}{B+\frac{1}{2}}\mu(x(k,q))$
+$\leqslant 2^{-N}\leqslant 2^{-q}.$
+The penultimate inequality above holds as $\mu(x(k,q))\leqslant\mu(z_{k})+2^{-1}\leqslant B+\frac{1}{2}$. Given that $\mu(z_{k}-Tz_{k})\leqslant 2^{-n}$, by (4.6) and the above we conclude that
+$\mu\bigg(x(k,q)-\sum_{i<N}x(k,q)(i)v_{i}\bigg)\leqslant 2^{-n}+(L+2)\cdot 2^{-q}.$
+Therefore $(\mu,x)\in AP_{n,L,N,v}$. The proof is complete. ∎

### v3

@@
 Recall that the map ${\delta}:{M}\to{\operatorname{Lip}_{0}(M)^{*}}$ given by $\langle{f},{\delta(x)}\rangle=f(x)$, $x\in M$, $f\in\operatorname{Lip}_{0}(M)$, is an isometric embedding, and that we can define the Lipschitz-free space (hereafter free space) $\mathcal{F}({M})$ over $M$ as the closed linear span of the image $\delta(M)\subseteq\operatorname{Lip}_{0}(M)^{*}$ (where $\delta(M\setminus\{0\})$ is linearly independent).
-For essential information on the fundamental properties of free spaces, we refer the reader to [godefroy:kalton:03, weaver:18]; in the latter they are known as Arens-Eells spaces.
+For essential information on the fundamental properties of free spaces, we refer the reader to [18, 31]; in the latter they are known as Arens-Eells spaces.
 Free spaces have close ties to metric geometry, optimal transport theory and the linear and non-linear theory of Banach spaces, and have been the subject of intense study by the functional analysis community in the last two decades.
@@
 In particular, it is not easy to find families of metric spaces whose corresponding free spaces are pairwise non-isomorphic.
-The first example of an uncountable family of separable metric spaces whose free spaces are pairwise non-isomorphic was given in [hajek:lancien:pernecka:16].
+The first example of an uncountable family of separable metric spaces whose free spaces are pairwise non-isomorphic was given in [21].
 These metric spaces are all homeomorphic to the Cantor set $2^{\omega}$.
-Two more uncountable families of this kind are given in [basset:24, basset:lancien:prochazka:25].
-In [basset:24], each metric space in the family is countable and complete, while in [basset:lancien:prochazka:25] each one is countable complete and discrete.
+Two more uncountable families of this kind are given in [4, 3].
+In [4], each metric space in the family is countable and complete, while in [3] each one is countable complete and discrete.
 The free spaces corresponding to the first family are non-isomorphic because the weak-fragmentability (or weak Szlenk) index of each one is distinct (and countable).
@@
 These results have consequences for universality.
-Recall that a metric space is purely 1-unrectifiable (p1u) if and only if it admits no Lipschitz isomorphic copy of any compact subset of $\mathbb{R}$ having strictly positive Lebesgue measure [kirchheim:94].
-By [aliaga:gartand:petitjean:prochazka:22], a complete separable metric space $M$ is p1u if and only if $\mathcal{F}({M})$ has the Radon-Nikodým property (RNP); equivalently the dentability index of $\mathcal{F}({M})$ is countable.
+Recall that a metric space is purely 1-unrectifiable (p1u) if and only if it admits no Lipschitz isomorphic copy of any compact subset of $\mathbb{R}$ having strictly positive Lebesgue measure [24].
+By [1], a complete separable metric space $M$ is p1u if and only if $\mathcal{F}({M})$ has the Radon-Nikodým property (RNP); equivalently the dentability index of $\mathcal{F}({M})$ is countable.
 Because the dentability index of each free space associated with the second family is distinct, the set of indices must be unbounded in $\omega_{1}$.
 Thus, if a complete separable metric space has a free space that is isomorphically universal for the class of free spaces over countable complete discrete metric spaces, it cannot be p1u.
-Consequently, if a complete separable metric space is Lipschitz universal for the class of countable complete discrete metric spaces, it cannot be p1u [basset:lancien:prochazka:25]*Corollary 4.7.
-In this paper we use a tool from descriptive set theory, specifically Bossard’s reduction argument (see [dodos:10]*p. 31), to build on this universality result.
+Consequently, if a complete separable metric space is Lipschitz universal for the class of countable complete discrete metric spaces, it cannot be p1u [3]*Corollary 4.7.
+In this paper we use a tool from descriptive set theory, specifically Bossard’s reduction argument (see [11]*p. 31), to build on this universality result.
 Theorem 1.1.
@@
 We say that $X$ has the bounded approximation property (BAP) if it has the $\lambda$-BAP for some $\lambda$, and the metric approximation property (MAP) if it has the $1$-BAP.
-The literature on free spaces and approximation properties is too abundant to be summarised in this short paper; we invite the interested reader to consult [godefroy:24] and the introduction of [smith:talimdjioski:23] and references therein for more information.
+The literature on free spaces and approximation properties is too abundant to be summarised in this short paper; we invite the interested reader to consult [20] and the introduction of [25] and references therein for more information.
 We only need to recall two results for our purposes.
-First, it follows immediately from [godefroy:kalton:03]*Theorem 3.1 that if a separable Banach space $X$ fails the AP then so does $\mathcal{F}({X})$.
-Second, [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8 implies that $\mathcal{F}({M})$ has the AP whenever $M$ is countable discrete and complete.
+First, it follows immediately from [18]*Theorem 3.1 that if a separable Banach space $X$ fails the AP then so does $\mathcal{F}({X})$.
+Second, [2]*Corollary 2.8 implies that $\mathcal{F}({M})$ has the AP whenever $M$ is countable discrete and complete.
 To the best of my knowledge, to date there has been no example of a complete separable p1u metric space whose free space has the AP but fails the BAP.
@@
 Again, this result is based on Bossard’s argument.
-Another crucial ingredient is the fact that, relative to the established descriptive set-theoretic frameworks for separable Banach spaces (see Section 2), the class of separable Banach spaces that satisfies the BAP is Borel (though being analytic would have sufficed) [ghawadrah:17].
+Another crucial ingredient is the fact that, relative to the established descriptive set-theoretic frameworks for separable Banach spaces (see Section 2), the class of separable Banach spaces that satisfies the BAP is Borel (though being analytic would have sufficed) [17].
 Theorem 1.3.
 There exist countable complete discrete metric spaces $M$ such that $\mathcal{F}({M})$ fails the BAP. Consequently, $\mathcal{F}({M})$ is not isomorphic to a dual space.
-This result can be contrasted with [aliaga:gartand:petitjean:prochazka:22]*Theorem 3.2 which states, among other things, that $\mathcal{F}({M})$ is isometric to a dual space whenever $M$ is a p1u metric space that is proper, i.e.
+This result can be contrasted with [1]*Theorem 3.2 which states, among other things, that $\mathcal{F}({M})$ is isometric to a dual space whenever $M$ is a p1u metric space that is proper, i.e.
 its closed balls are compact.
-It can also be regarded in the light of Kalton’s problem of whether $\mathcal{F}({M})$ has the BAP whenever $M$ is uniformly discrete; see [kalton:04]*p.
+It can also be regarded in the light of Kalton’s problem of whether $\mathcal{F}({M})$ has the BAP whenever $M$ is uniformly discrete; see [22]*p.
 185 and e.g.
-[godefroy:24]*Problem 4.1 for more context.
-In this vein, there exists a uniformly discrete space whose free space is not isometric to a dual space (though this particular example has the MAP) [garcia-lirola:petitjean:prochazka:ruedazoca:18]*Example 5.8.
+[20]*Problem 4.1 for more context.
+In this vein, there exists a uniformly discrete space whose free space is not isometric to a dual space (though this particular example has the MAP) [16]*Example 5.8.
 We prove the results above using Lemma 3.10, which is one of our two main tools.
 The other main tool, Lemma 3.9, will be needed for our final main result, Theorem 4.1, which shows that the classes of separable Banach spaces and separable free spaces having the AP are, relative to the appropriate descriptive set-theoretic interpretations, complete coanalytic sets.
-(We remark that this is not the first time Polish spaces of metrics have been used to study free spaces and approximation properties; see [talimdjioski:23, smith:talimdjioski:24, talimdjioski:25].)
+(We remark that this is not the first time Polish spaces of metrics have been used to study free spaces and approximation properties; see [27, 26, 28].)
 We prove these theorems in Sections 3 and 4, and make some related observations along the way.
 The application of descriptive set theory to Banach space theory can be traced to some seminal papers of Bourgain, e.g.
-[bourgain:80].
-It was formalised later by Bossard in another seminal paper [bossard:02] and has since developed into a rich and beautiful subfield that boasts many powerful techniques and results, see e.g.
-[dodos:10] and references therein, and [cuth:doucha:kurka:22b] and related papers for more recent developments.
+[7].
+It was formalised later by Bossard in another seminal paper [6] and has since developed into a rich and beautiful subfield that boasts many powerful techniques and results, see e.g.
+[11] and references therein, and [9] and related papers for more recent developments.
 I speculate that techniques from descriptive set theory could be used to prove more results about free spaces in the future.
-I heard about the results in [basset:lancien:prochazka:25] in a talk given by E.
+I heard about the results in [3] in a talk given by E.
 Basset at the Banff International Research Station workshop on Functional and Metric Analysis and their Interactions, held at the Institute of Mathematics, University of Granada, Spain, in May 2025.
-It occurred to me that it might be possible to use descriptive set theory, specifically Bossard’s reduction argument (see [dodos:10]*p.
+It occurred to me that it might be possible to use descriptive set theory, specifically Bossard’s reduction argument (see [11]*p.
 31), to build on their universality results.
@@
 Rueda Zoca for several remarks which helped me when I was preparing this paper.
-Finally, I should stress that I have left out most of the results in [basset:24, basset:lancien:prochazka:25] from this introduction purely in the interests of brevity, and I recommend that the interested reader should consult these papers in full.
+Finally, I should stress that I have left out most of the results in [4, 3] from this introduction purely in the interests of brevity, and I recommend that the interested reader should consult these papers in full.
 2. Notation and preliminaries
@@
 We also use these symbols in the context of metric spaces $M$ and $N$, with “Lipschitz” substituted for “linear”.
-By [weaver:18]*Theorems 3.6 and 3.7, if $M\hookrightarrow N$ (in the Lipschitz sense) then $\mathcal{F}({M})\hookrightarrow\mathcal{F}({N})$ (in the linear sense).
+By [31]*Theorems 3.6 and 3.7, if $M\hookrightarrow N$ (in the Lipschitz sense) then $\mathcal{F}({M})\hookrightarrow\mathcal{F}({N})$ (in the linear sense).
 We say that the Banach space $X$ is isomorphically universal for a given class $\mathcal{C}$ of Banach spaces if $Y\hookrightarrow X$ whenever $Y\in\mathcal{C}$; Lipschitz universality of metric spaces is defined analogously.
-For concepts pertaining to descriptive set theory, we follow [kechris:95].
+For concepts pertaining to descriptive set theory, we follow [23].
 The sets of natural numbers, countable ordinals and infinite sequences of natural numbers are denoted $\omega$, $\omega_{1}$ and $\omega^{\omega}$, respectively.
@@
 We shall consider the Polish space $0,\infty)^{\omega\times\omega}$ endowed with the pointwise topology, and the set $\mathcal{R}\subseteq[0,\infty)^{\omega\times\omega}$ of metrics on $\omega$.
-This space was introduced by Vershik in [vershik:98] and features again in e.g.
-[vershik:04] and [clemens:12] (in these works $\mathcal{R}$ is defined instead as the set of all pseudo-metrics on $\omega$, that is, functions $d\in[0,\infty)^{\omega\times\omega}$ that are metrics except that it is possible to have distinct $i,j\in\omega$ satisfying $d(i,j)=0$).
+This space was introduced by Vershik in [29] and features again in e.g.
+[30] and [8] (in these works $\mathcal{R}$ is defined instead as the set of all pseudo-metrics on $\omega$, that is, functions $d\in[0,\infty)^{\omega\times\omega}$ that are metrics except that it is possible to have distinct $i,j\in\omega$ satisfying $d(i,j)=0$).
 It is easily seen that $\mathcal{R}$ is a $G_{\delta}$ in $0,\infty)^{\omega\times\omega}$ and is thus Polish as well.
@@
 Bossard coded the class of separable infinite-dimensional Banach spaces as a $\mathbf{\Delta}^{1}_{1}$ subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure.
-Following a programme initiated in [godefroy:saint-raymond:18], the authors of [cuth:doucha:kurka:22a, cuth:doucha:kurka:22b] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$.
+Following a programme initiated in [19], the authors of [10, 9] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$.
 First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ whose canonical extension to $c_{00}$ is a norm.
@@
 While not directly relevant to this paper, another advantage of using $\mathcal{B}$ is that Baire category arguments can be applied, e.g.
-the set of $\mu\in\mathcal{B}$ such that $X_{\mu}\equiv\mathbb{G}$, where $\mathbb{G}$ is the Gurariĭ space, is a dense $G_{\delta}$ [cuth:doucha:kurka:22b]*Theorem 4.1, hence $\mathbb{G}$ can be thought of as the generic separable Banach space.
-According to [cuth:doucha:kurka:22a]*Theorem 8, there is a Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ such that $\Theta(\mu)\equiv X_{\mu}$ for all $\mu\in\mathcal{B}$.
-It follows easily from this and [bossard:02]*Theorem 2.3 (ii) that, given any separable Banach space $X$, the set $\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow}:=\left\{{\mu\in\mathcal{B}}\;:\;\,{X_{\mu}\hookrightarrow X}\right\}$ is $\mathbf{\Sigma}^{1}_{1}$.
+the set of $\mu\in\mathcal{B}$ such that $X_{\mu}\equiv\mathbb{G}$, where $\mathbb{G}$ is the Gurariĭ space, is a dense $G_{\delta}$ [9]*Theorem 4.1, hence $\mathbb{G}$ can be thought of as the generic separable Banach space.
+According to [10]*Theorem 8, there is a Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ such that $\Theta(\mu)\equiv X_{\mu}$ for all $\mu\in\mathcal{B}$.
+It follows easily from this and [6]*Theorem 2.3 (ii) that, given any separable Banach space $X$, the set $\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow}:=\left\{{\mu\in\mathcal{B}}\;:\;\,{X_{\mu}\hookrightarrow X}\right\}$ is $\mathbf{\Sigma}^{1}_{1}$.
 Concerning Polish versus Effros-Borel structures, we should mention that it is possible to code the class of complete infinite separable metric spaces using the set of closed infinite subsets of the Urysohn universal separable metric space $\mathbb{U}$, again equipped with the Effros-Borel structure, see e.g.
-[gao:09, gao:kechris:03].
+[15, 14].
 However we shall work with Vershik’s space $\mathcal{R}$ for the reasons given above.
-Also, analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [vershik:04]*Theorems 1 and 3.
-Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [fonf:wojtaszczyk:08], could be regarded as the generic separable free space.
-A immediate consequence of this and [fonf:wojtaszczyk:08]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual.
+Also, analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [30]*Theorems 1 and 3.
+Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [13], could be regarded as the generic separable free space.
+A immediate consequence of this and [13]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual.
 This observation complements Theorem 4.1.
@@
 Then $\Xi(d^{\prime})(v_{\ell})=\Xi(d)(v_{\ell})$ for $\ell=1,\ldots,k$.
-This holds because, given $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$, by McShane’s extension theorem [weaver:18]*Theorem 1.33 there exists $g\in B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$, and vice-versa.
+This holds because, given $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$, by McShane’s extension theorem [31]*Theorem 1.33 there exists $g\in B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$, and vice-versa.
 Therefore $\Xi$ is continuous.
@@
 Proof.
-Evidently $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\Xi^{-1}(\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$. Similarly to [clemens:12]*Lemma 4, the set
+Evidently $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\Xi^{-1}(\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$. Similarly to [8]*Lemma 4, the set
 $\left\{{(d,d^{\prime})\in\mathcal{R}\times\mathcal{R}}\;:\;\,{M_{d}\hookrightarrow M_{d^{\prime}}}\right\}$
@@
 Proof.
-Using the Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ and [bossard:02]*Corollary 3.3, the set of $\mu\in\mathcal{B}$ such that $X_{\mu}$ has the RNP is $\mathbf{\Pi}^{1}_{1}$.
-Now apply Proposition 3.1 and [aliaga:gartand:petitjean:prochazka:22]*Theorem 4.6.
+Using the Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ and [6]*Corollary 3.3, the set of $\mu\in\mathcal{B}$ such that $X_{\mu}$ has the RNP is $\mathbf{\Pi}^{1}_{1}$.
+Now apply Proposition 3.1 and [1]*Theorem 4.6.
 ∎
@@
 $P=\left\{{E\in 2^{\omega\times\omega}}\;:\;\,{\left\{{j\in\omega}\;:\;\,{(i,j)\in E}\right\}\text{ is finite for all }i\in\omega}\right\}.$
-According to [kechris:95]*Exercise 23.1, $P$ is $\mathbf{\Pi}^{0}_{3}$-complete.
+According to [23]*Exercise 23.1, $P$ is $\mathbf{\Pi}^{0}_{3}$-complete.
 Now set $\Gamma=\omega\times\omega$, $M=\ell_{1}(\Gamma\cup\{\infty\})$, $A=\{0\}$ and $\phi_{0}=\{(0,0)\}$.
@@
 ∎
-The final lemma below uses Bossard’s reduction argument [bossard:02]*Theorem 3.2; Theorems 1.11.3 emerge as relatively straightforward consequences of it.
+The final lemma below uses Bossard’s reduction argument [6]*Theorem 3.2; Theorems 1.11.3 emerge as relatively straightforward consequences of it.
 Lemma 3.10.
@@
 Let $d^{\prime}\in\mathcal{R}$ and $d\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$.
-By [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})\hookrightarrow\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
+By [31]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})\hookrightarrow\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
 Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
 By Lemma 3.10, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$.
-As $Y\hookrightarrow\mathcal{F}({c_{0}})$ whenever $Y$ is a separable Banach space ([dutrieux:ferenczi:06]*p.
+As $Y\hookrightarrow\mathcal{F}({c_{0}})$ whenever $Y$ is a separable Banach space ([12]*p.
 1042) and $c_{0}\equiv M_{d}$ for some $d\in\mathcal{R}$, it follows that $Y\hookrightarrow X$ for all such $Y$.
@@
 $\mathbb{U}$ or $c_{0}$ (the latter being Lipschitz universal by Aharoni’s theorem, see e.g.
-[benyamini:lindenstrauss:00]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal.
+[5]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal.
 ∎
 Proof of Theorem 1.3.
-Using [ghawadrah:17], [cuth:doucha:kurka:22a]*Theorem 8 and Proposition 3.1, the set
+Using [17], [10]*Theorem 8 and Proposition 3.1, the set
 […]
@@
 Let $d^{\prime}\in\mathcal{R}$ and $d\in\mathcal{R}_{\mathrm{BAP}}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$.
-Again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\mathcal{R}_{\mathrm{BAP}}$.
+Again by [31]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\mathcal{R}_{\mathrm{BAP}}$.
 For a contradiction assume $\mathcal{R}_{\mathrm{dis}}\subseteq\mathcal{R}_{\mathrm{BAP}}$.
 Then $\mathcal{R}_{\mathrm{BAP}}=\mathcal{R}$ by Lemma 3.10.
-However, this is false because we can pick $d\in\mathcal{R}$ such that $M_{d}$ is isometric to a separable Banach space that fails the AP; as is pointed out in the Introduction, $d\notin\mathcal{R}_{\mathrm{BAP}}$ by [godefroy:kalton:03]*Theorem 3.1.
+However, this is false because we can pick $d\in\mathcal{R}$ such that $M_{d}$ is isometric to a separable Banach space that fails the AP; as is pointed out in the Introduction, $d\notin\mathcal{R}_{\mathrm{BAP}}$ by [18]*Theorem 3.1.
 Now pick $d\in\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$.
-Again, as stated in the Introduction, $\mathcal{F}({M_{d}})$ has the AP by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8.
+Again, as stated in the Introduction, $\mathcal{F}({M_{d}})$ has the AP by [2]*Corollary 2.8.
 That $\mathcal{F}({M_{d}})$ cannot be isomorphic to a dual space is a simple consequence of the classical result of Grothendieck which asserts that a separable Banach space has the MAP whenever it has the AP and is isometric to a dual space.
@@
   1. (1)
-    Given metric spaces $M$ and $N$, let us write $M\stackrel{{\scriptstyle u}}{{\hookrightarrow}}N$ if $M$ is uniformly homeomorphic to a subspace of $N$. Similarly to [clemens:12]*Lemma 4, the set
+    Given metric spaces $M$ and $N$, let us write $M\stackrel{{\scriptstyle u}}{{\hookrightarrow}}N$ if $M$ is uniformly homeomorphic to a subspace of $N$. Similarly to [8]*Lemma 4, the set
     $\{(d,d^{\prime})\in\mathcal{R}\times\mathcal{R}\,:\,M_{d}\stackrel{{\scriptstyle u}}{{\hookrightarrow}}M_{d^{\prime}}\}$
@@
 No changes to the proofs above are needed.
-Given $d\in\mathcal{R}_{\mathrm{p1u}}$, we have $\langle\mathcal{F}({M_{d}})\rangle^{\mathcal{R}}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{p1u}}$ by [aliaga:gartand:petitjean:prochazka:22]*Theorem 4.6.
+Given $d\in\mathcal{R}_{\mathrm{p1u}}$, we have $\langle\mathcal{F}({M_{d}})\rangle^{\mathcal{R}}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{p1u}}$ by [1]*Theorem 4.6.
 Therefore, by arguing as above, there is a sequence $(d_{\alpha})_{\alpha<\omega_{1}}$ of elements of $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ such that $\mathcal{F}({M_{d_{\beta}}})\not\hookrightarrow\mathcal{F}({M_{d_{\alpha}}})$ whenever $\alpha<\beta<\omega_{1}$.
@@
 Apply Lemma 3.9 to a separable Banach space $M$ that fails the AP.
-If $T\in\mathbf{WF}$ then $\mathcal{F}({M_{\Delta(T)}})$ has the AP again by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8.
+If $T\in\mathbf{WF}$ then $\mathcal{F}({M_{\Delta(T)}})$ has the AP again by [2]*Corollary 2.8.
 If $T\in\mathbf{IF}$ then $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$.
-Once again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{\Delta(T)}})$.
+Once again by [31]*Theorems 3.6 and 3.7, $\mathcal{F}({M})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{\Delta(T)}})$.
 Since $\mathcal{F}({M})$ fails the AP, so must $\mathcal{F}({M_{\Delta(T)}})$.
@@
 This section is solely concerned with determining the complexity of two classes of Banach spaces having the AP.
-The proof of the result below is an adaptation of an argument from [ghawadrah:17].
+The proof of the result below is an adaptation of an argument from [17].
 As above, we frame the result in terms of the Polish space $\mathcal{B}$ instead of the Effros-Borel space $\mathcal{SB}$.
@@
 Hence we can define a bounded $\mathbb{Q}$-linear operator ${T}:{V}\to{\operatorname{span}(e_{i})_{i<R}}$ by setting $Te_{i}=w_{i}$, $i\in\omega$, and then extending to finite $\mathbb{Q}$-linear combinations of the $e_{i}$.
-By equicontinuity we extend it again to a bounded linear operator on $X_{\mu}$ satisfying $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$ and $\mu(T)\leqslant L$.
+By uniform continuity we extend it again to a bounded linear operator on $X_{\mu}$ satisfying $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$ and $\mu(T)\leqslant L$.
 By taking a further subsequence of natural numbers, also labelled $(N_{\ell})$, we can ensure that

### v4

@@
 Following a programme initiated in [18], the authors of [9, 8] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$.
-First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ whose canonical extension to $c_{00}$ is a norm.
+First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ which can be extended (uniquely) to a norm on $c_{00}$.
 Endowed with the pointwise topology, $\mathcal{B}$ is a $G_{\delta}$ in $0,\infty)^{V}$ and is thus Polish.
 Every separable and infinite-dimensional Banach space $(X,\left\|\cdot\right\|)$ has representatives in $\mathcal{B}$.
-Given a dense linearly independent sequence $(x_{i})\subseteq X$, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$.
+Given a linearly independent sequence $(x_{i})\subseteq X$ having dense span, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$.
 Then $X\equiv X_{\mu}$, where $X_{\mu}$ is the completion of $(V,\mu)$; conversely, every such completion is a separable and infinite-dimensional Banach space.
@@
 Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [12], could be regarded as the generic separable free space.
-A immediate consequence of this and [13]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual.
+An immediate consequence of this and [12]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual.
 This observation complements Theorem 4.1.
@@
 we quickly conclude that $X_{\Xi(d)}\equiv\mathcal{F}({M_{d}})$.
-The map $\Xi$ is continuous.
-Indeed, let $d\in\mathcal{R}$ and $v_{1},\ldots,v_{k}\in V$.
-There exists $N\in\omega$ such that $v_{\ell}(i)=0$ whenever $\ell=1,\ldots,k$ and $i\geqslant N$.
-Consider any $d^{\prime}\in\mathcal{R}$ such that $d^{\prime}(i,j)=d(i,j)$ whenever $i,j\leqslant N$.
-Then $\Xi(d^{\prime})(v_{\ell})=\Xi(d)(v_{\ell})$ for $\ell=1,\ldots,k$.
-This holds because, given $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$, by McShane’s extension theorem [31]*Theorem 1.33 there exists $g\in B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$, and vice-versa.
-Therefore $\Xi$ is continuous.
-∎
+It remains to show that $\Xi$ is continuous.
+Let $d\in\mathcal{R}$, $v_{1},\ldots,v_{k}\in V$ and $\varepsilon\in(0,1)$.
+To establish the continuity of $\Xi$, we shall find $N\in\omega$ and $\delta>0$ such that
+$|\Xi(d^{\prime})(v_{\ell})-\Xi(d)(v_{\ell})|<\varepsilon,\quad 1\leqslant\ell\leqslant k,$
+whenever $d^{\prime}\in\mathcal{R}$ satisfies $|d^{\prime}(i,j)-d(i,j)|<\delta$ for all $i,j\leqslant N$.
+To begin, we pick $N\in\omega$ such that $v_{\ell}(i)=0$ whenever $1\leqslant\ell\leqslant k$ and $i\geqslant N$. Then set
+$r=\min\left\{{d(i,j)}\;:\;\,{i<j\leqslant N}\right\}>0,\quad L=1+\max\left\{{\Xi(d)(v_{\ell})}\;:\;\,{1\leqslant\ell\leqslant k}\right\},$
+$\alpha=\varepsilon/(L+1)$ and $\delta=\alpha r$.
+Now consider any $d^{\prime}\in\mathcal{R}$ such that $|d^{\prime}(i,j)-d(i,j)|<\delta$ whenever $i,j\leqslant N$. Let $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$. Given that
+$(1-\alpha)d(i,j)\leqslant d(i,j)-\delta<d^{\prime}(i,j)<d(i,j)+\delta\leqslant(1+\alpha)d(i,j),\quad i<j\leqslant N,$ (3.1)
+the restriction of $f$ to $N+1$ is $(1+\alpha)$-$d$-Lipschitz.
+By McShane’s extension theorem [30]*Theorem 1.33, there exists $g\in(1+\alpha)B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$.
+It follows that
+$\sum_{i\in\omega}v_{\ell}(i)f(i+1)=\sum_{i<N}v_{\ell}(i)f(i+1)$ $=\sum_{i<N}v_{\ell}(i)g(i+1)$
+$=\sum_{i\in\omega}v_{\ell}(i)g(i+1)\leqslant(1+\alpha)\Xi(d)(v_{\ell}),\quad 1\leqslant\ell\leqslant k.$
+As these inequalities hold for all such $f$, we obtain
+$\Xi(d^{\prime})(v_{\ell})\leqslant(1+\alpha)\Xi(d)(v_{\ell})<\Xi(d)(v_{\ell})+\varepsilon<L,\quad 1\leqslant\ell\leqslant k.$
+Now we swap $d$ and $d^{\prime}$ and largely repeat the above.
+Given $f\in B_{\operatorname{Lip}_{0}(M_{d})}$, by (3.1) and McShane’s theorem, we can find $g\in(1-\alpha)^{-1}B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$.
+It follows that
+$\Xi(d)(v_{\ell})\leqslant\frac{1}{1-\alpha}\Xi(d^{\prime})(v_{\ell})=\bigg(1+\frac{\varepsilon}{L}\bigg)\Xi(d^{\prime})(v_{\ell})<\Xi(d^{\prime})(v_{\ell})+\varepsilon,\quad 1\leqslant\ell\leqslant k.$
+This completes the proof. ∎
 Proposition 3.2.
@@
 Lemma 3.7.
-Let $M=(M,\theta)$ be a separable metric space, let $(x_{i})_{i\in\omega}$ be a dense sequence in $M$ whose elements repeat infinitely often, let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.1) and define a second metric $\eta$ on $\omega^{<\omega}$ by
+Let $M=(M,\theta)$ be a complete separable metric space, let $(x_{i})_{i\in\omega}$ be a dense sequence in $M$ whose elements repeat infinitely often, let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.2) and define a second metric $\eta$ on $\omega^{<\omega}$ by
 $\eta(s,t)=\rho(s,t)+\theta(x_{|s|},x_{|t|}).$
@@
 Lemma 3.8.
-Let $\eta$ be the metric from Lemma 3.7, where $(M,\theta)$ is arbitrary and $\rho$ is as above.
+Let $\eta$ be the metric from Lemma 3.7, where $(M,\theta)$ is an arbitrary separable metric space and $\rho$ is as above.
 Then $(\omega^{<\omega},\eta)$ is discrete, and in addition $(T,\eta)$ is complete whenever $T\in\mathbf{WF}$.
@@
 Lemma 3.9.
-Let $M=(M,\theta)$ be a separable metric space.
+Let $M=(M,\theta)$ be a complete separable metric space.
 There exists a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ whenever $T\in\mathbf{WF}$ and $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$ whenever $T\in\mathbf{IF}$.
@@
 Let $(\omega^{<\omega},\eta)$ be the metric space obtained from Lemma 3.7 applied to $M$.
-Let $W=\omega^{<\omega}\cup(\omega^{<\omega}\cup\{0\})$.
+Let $W=\omega^{<\omega}\cup(\omega^{<\omega}\times\{0\})$.
 We extend $\eta$ to a metric on $W$, again labelled $\eta$, by setting
@@
 By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$.
-Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
+Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$.
 By Lemma 3.10, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$.
-By considering e.g.
-$\mathbb{U}$ or $c_{0}$ (the latter being Lipschitz universal by Aharoni’s theorem, see e.g.
-[5]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal.
 ∎